### v1

# PoIFusion: Multi-Modal 3D Object Detection via Fusion at Points of Interest

Abstract
In this work, we present PoIFusion, a simple yet effective multi-modal 3D object detection framework to fuse the information of RGB images and LiDAR point clouds at the point of interest (abbreviated as PoI). Technically, our PoIFusion follows the paradigm of query-based object detection, formulating object queries as dynamic 3D boxes. The PoIs are adaptively generated from each query box on the fly, serving as the keypoints to represent a 3D object and play the role of basic units in multi-modal fusion. Specifically, we project PoIs into the view of each modality to sample the corresponding feature and integrate the multi-modal features at each PoI through a dynamic fusion block. Furthermore, the features of PoIs derived from the same query box are aggregated together to update the query feature. Our approach prevents information loss caused by view transformation and eliminates the computation-intensive global attention, making the multi-modal 3D object detector more applicable. We conducted extensive experiments on the nuScenes dataset to evaluate our approach. Remarkably, our PoIFusion achieves 74.9% NDS and 73.4% mAP, setting a state-of-the-art record on the multi-modal 3D object detection benchmark. Codes will be made available via https://djiajunustc.github.io/projects/poifusion.
1 Introduction
Autonomous vehicles are usually equipped with an array of sensors to facilitate safe driving, among which cameras and LiDARs are the most popular. These two sensors are complementary to each other: cameras provide rich textual and color information, while LiDAR sensors supply precise spatial measurements. The effective camera-LiDAR data fusion is widely regarded as a promising direction to achieve high-quality 3D object detection [2, 34, 42, 64], which has attracted a surge of research interest in the community.
Figure 1: A comparison of (a) fusion with unified view, (b) fusion with global attention, and (c) our PoIFusion. “VT” and “PC” indicate view transformation and point clouds, respectively.
The fundamental challenge of camera-LiDAR data fusion arises from the discrepancy of their representation space (i.e., 2D perspective view versus 3D space). To ameliorate this challenge, one common solution is to transform the image and point cloud representations into a unified bird-eye view [42, 34, 22], as depicted in Figure 1(a), or into 3D space [29]. Another recent approach [2, 69], as shown in Figure 1(b), keeps the representation in its original view and abstracts multi-modal features into object queries [6, 40] with the global attention mechanism [56].
However, both of these approaches have inherent issues. In the unified-view approach, the core component, i.e., view transformation, is often based on monocular depth estimation to lift the 2D image into 3D. However, depth estimation is an error-prone task, with errors having a deleterious effect on any downstream tasks (such as object recognition). Moreover, the direct grid-to-grid fusion [34, 42] loses a significant portion of the original representational strengths which comprises modal-specific information [75].
The second query-based approach can avoid feature ambiguity and information loss by keeping the original view of the feature representation. However, the adoption of global attention to integrating multi-modal features incurs high computation and memory overhead. For instance, the state-of-the-art algorithm [69], which follows the query-based approach of fusion with global attention, relies on the well-optimized Flash Attention operator [15] to cut down the time and memory consumption. The high overhead has become an obstacle that hinders the wide application of the algorithm. Furthermore, it is difficult to extract the object-relevant feature with the global attention mechanism [84, 21], especially when it comes to such a large 3D space like the autonomous driving scenario.
The benefit of the query-based approach inspires us to preserve the original view of features from each modality, while the above two issues motivate our exploration of an alternative paradigm to replace dense feature interaction of the global attention with sparse point projection and feature sampling, as illustrated in Figure 1(c).
As such, we propose a new query-based multi-modal 3D object detection framework that initializes object queries as learnable 3D boxes and dynamically integrates multi-modal features at representative points derived from each object query. In this manuscript, these representative points are referred to as Points of Interest (PoIs), and our framework is named PoIFusion. Intuitively, a naive way to represent a 3D query box with points is to use the center point [8, 69, 40]. However, simply representing a 3D box with its center point totally ignores the geometric properties, such as the size and rotation angle. For supplementary, an improved design is to involve the corners [49]. Nevertheless, sampling multi-modal features according to the projected location of center and corner points also incurs the problem of feature misalignment, since the projection of a 3D box may not be a tight bounding box onto the image view. The issue remains even if a box is well-located in the 3D space [5], not to mention that the query box is not guaranteed to be accurately localized. To ameliorate this problem, our PoIs are adaptively generated from the center and corner points, with box-level and point-level transformation parameters online predicted according to the query feature. In our design, a single PoI serves as the basic unit for fine-grained multi-modal feature fusion, and the ensemble of PoIs derived from the same object query represents the regional feature in a flexible way.
Once PoIs are generated, the features from different modalities can be easily obtained by projecting PoIs onto the corresponding view, followed by bilinear interpolation according to the projected location. Moreover, because different modalities contribute differently to each object query, a dynamic fusion block is engaged in our design. Particularly, in our dynamic fusion block, we first generate the parameters of the fusion layer on the fly, and then integrate the sampled multi-modal feature at each PoI. The proposed adaptive PoIs, together with the dynamic fusion block, enable our PoIFusion to efficiently sample the object-relevant multi-modal features and make the best of modal-specific information, thus improving 3D object detection.
We benchmark the proposed PoIFusion framework on the nuScene dataset and conduct extensive experiments to validate our designs. Notably, our method sets a new record on this challenging dataset, achieving 74.9% NDS and 73.4% mAP on the test set without bells and whistles.
In summary, our contributions are three-fold:
  • We propose a novel PoIFusion framework for multi-modal 3D object detection that preserves modality-specific representation spaces while efficiently extracting and fusing features through sparse interactions.
  • We present the design of fusion at points of interest, conveying an elegant view that the entity involved in the fusion module can be very flexible.
  • We conduct extensive experiments to validate the effectiveness of our method, demonstrating its potential to serve as a strong baseline for future investigation.
2 Related Work
LiDAR-Based 3D Object Detection.
LiDAR sensors capture 3D point clouds, which provide accurate spatial information that can be important for 3D object detection. Broadly, 3D object detection algorithms operated on point clouds can be categorized into two groups: point-based and voxel-based ones. Point-based 3D object detection algorithms [51, 74, 73] make direct use of the precise coordinates of points, progressively sampling keypoints and extracting local information with set abstraction operators [46, 47, 52, 53] that aggregate information form a point and its neighbours. In contrast, voxel-based 3D object detection algorithms [80, 72, 28, 16, 76, 50, 20, 17] first “voxelize” the point cloud by binning points into a regulare grid. The voxel representation enables straightforward feature extraction using standard (or sparse) convolutions [70, 9, 10] or sparse voxel Transformers [19, 59, 81, 27, 12, 18, 44].
Figure 2: An overview of our proposed PoIFusion framework, which is mainly composed of (a) an image feature module, (b) a point cloud feature module, and (c) a multi-modal object decoder. In our PoIFusion, we independently extract the feature of each modality and keep the original representation view (i.e., image feature in the perspective view, and point cloud feature in the bird-eye view). The multi-modal feature maps are taken as the input of our multi-modal object decoder. The object decoder is iteratively applied 6 times to integrate the multi-modal feature sampled by generated points of interest (PoIs) and refine the object queries.
Camera-Based 3D Object Detection.
3D object detection is one of the oldest Computer Vision problems and a variety of approaches have been proposed over many years (we do not survey these here). For the purposes of 3D object detection from a moving car a popular approach, and one that can be readily fused with point-cloud estimates, is to measure detection success in the Birds-eye View (BEV) space. Earlier attempts along these lines – e.g., geometric uncertainty [43] and pseudo-LiDAR representation [62] – mainly focus on monocular 3D object detection [61, 48, 3, 14, 79]. Autonomous driving vehicles are typically equipped with multiple cameras, providing perception information over the full 360 degrees. To leverage the relationship between multi-view images, BEV-based and query-based algorithms have been explored. BEV-based 3D object detection [25, 33, 32, 71] explicitly performs view transformation to unify multi-view images into bird-eye-view representation. Query-based 3D object detection [40, 63, 38, 65, 54, 68] follows the pipeline of DETR [6], capitalizing on learnable object queries to extract multi-view information without view transformation.
Multi-Modal 3D Object Detection.
Although the exploration of each individual modality has made encouraging progress, the accuracy and robustness of detection algorithms are still insufficient for safe driving. The fact that images and point clouds are naturally complementary to each other (i.e., rich semantic information versus precise spatial information) has motivated further exploration in multi-modal 3D object detection [60, 2, 8, 22, 67, 31, 26]. In the early stage, the multi-modal approaches utilize point clouds as the principal component, introducing image features at the point level [58, 57, 77, 13] or proposal level [83, 7, 78] to enhance the features of the point clouds. Recently, inspired by multi-view 3D object detection, a series of works [34, 42, 29] propose that unifying the representation space with explicit view transformation [45] facilitates multi-modal fusion. Another group of methods [30, 2, 69, 75] leverages the attention mechanism in Transformer architecture [56] to perform multi-modal fusion in a sequential [2] or a parallel [69] manner. A recent work, ObjectFusion [5], fuses multi-modal features in a two-stage pipeline. It first generates region proposals with the image-augmented BEV features [2], and then extracts the object-centric feature [24, 16] from the voxel, image, and BEV space for further fusion and refinement. In this work, the proposed PoIFusion is also object-centric. However, in contrast to integrating the corresponding region-level feature from multiple modalities, we utilize the PoIs as the basic units to perform multi-modal fusion, which is adaptively derived from the query box - this element stands as the cornerstone of our approach, fundamentally improving the efficacy and flexibility of our framework.
3 Our Approach
In this section, we present the details of our PoIFusion. In Section 3.1, we provide a comprehensive overview of our framework. Then, in Section 3.2, we explain how to generate PoIs based on the object query. Subsequently, in Section 3.3, we introduce the process of multi-modal feature sampling. After that, we elaborate on our dynamic fusion block design in Section 3.4. Finally, in Section 3.5, we detail our prediction head and the training objective.
3.1 Overview
As illustrated in Figure 2, our PoIFusion consists of three main components: (a) an image feature module, (b) a point cloud feature module, and (c) a multi-modal object decoder. Given multi-view images and point clouds, perspective-view image feature maps $\bm{F}_{I}$ and bird-eye-view (BEV) point cloud feature maps $\bm{F}_{P}$ are independently extracted with the camera encoder and the LiDAR encoder. The camera encoder is composed of an image backbone network and an FPN [35] to obtain multi-scale features. The LiDAR encoder exploits the VoxelNet [70, 80] and a BEV backbone network, following the common practice of the voxel-based paradigm. Then, the multi-modal object decoder, which works as the core component in our method, is iteratively applied 6 times, progressively integrating the multi-modal features and refining the detection boxes.
In our multi-modal decoder, each object query $\bm{Q}$ is formulated as an adaptive 3D box $\bm{b}_{t}$, associating with a feature vector $\bm{v}_{t}$ ($t$ indicates the iteration time step). In each iteration, the object queries are first fed into a self-attention layer to capture the relationships and dependencies between different objects. Here, we follow [38] to exploit a distance-biased self-attention layer in our decoder. Subsequently, for each object query, the box-wise transformation parameters and point-wise shift parameters are generated based on the query feature $\bm{v}_{t}$, and applied on the query box $\bm{b}_{t}$ to obtain a set of PoIs $\bm{P}_{t}=\{P^{i}\}$. To perform multi-modal fusion, a three-step process is undertaken for PoIs. Firstly, each PoI is projected onto both the perspective view and the bird-eye view, establishing the correspondence between the PoI and the multi-modal feature maps. Secondly, the image feature $\bm{f}_{I}$ and point cloud feature $\bm{f}_{P}$ of a PoI are sampled through bilinear interpolation, regarding the projected location on the corresponding view. Thirdly, the extracted multi-modal features at each PoI are dynamically integrated using a dynamic fusion block, followed by a feedforward network (FFN). Finally, a classification head and a regression head are applied to obtain the detection boxes, and the query feature and query box are updated, respectively.
3.2 PoI Generation
In this section, we elaborate on how to generate points of interest (PoIs), which serve as the basic units for multi-modal feature fusion in our PoIFusion framework.
[…]
$x_{c}^{\prime}$ $=x_{c}+t_{x},$ $w^{\prime}$ $=w\cdot\exp(t_{w}),$ (1)
$y_{c}^{\prime}$ $=y_{c}+t_{y},$ $l^{\prime}$ $=l\cdot\exp(t_{l}),$
$z_{c}^{\prime}$ $=z_{c}+t_{z},$ $h^{\prime}$ $=h\cdot\exp(t_{h}),$
$\sin\theta^{\prime}$ $=\sin\theta+t_{\sin},$ $\cos\theta^{\prime}$ $=\cos\theta+t_{\cos}.$
Once the box transformation has been applied, the center point and 8 corner points of $\bm{b^{\prime}}$ are collected together as the anchor points $\{A^{i}\}_{i=0}^{8}$. Finally, the point-wise shift is independently conducted on each anchor point $A^{i}$ to produce a PoI $P^{i}$ ($P^{i}=A^{i}+\Delta_{P}^{i}$). It’s worth noting that one object query corresponds to a set of PoIs derived from the center and corner points of the query box.
3.3 Feature Sampling
In our approach, we assemble multi-modal features by establishing the correspondence between a PoI and multi-modal feature maps via projection, followed by sampling features utilizing bilinear interpolation.
Let us denote the location of a 3D PoI $P^{i}$ as [$x^{i}$, $y^{i}$, $z^{i}$]. To sample the point cloud feature, $P^{i}$ is projected onto the bird-eye view (BEV). Since we exploit the voxel representation for point clouds, the location of the projected point on the BEV point cloud feature $F_{P}$ is computed as:
{$m^{i}&=\frac{x^{i}-X_{\text{min}}}{(X_{\text{max}}-X_{\text{min}})\times d},\\ n^{j}&=\frac{y^{i}-Y_{\text{min}}}{(Y_{\text{max}}-Y_{\text{min}})\times d},\end{aligned}\right.$ (2)
where [$m^{i}$, $n^{i}$] is the projected location, [$X_{\text{min}}$, $Y_{\text{min}}$, $X_{\text{max}}$, $Y_{\text{max}}$] is the point cloud range in the BEV, and $d=8$ is the downsampling scalar of point cloud feature maps. The point cloud feature sampled at [$m^{i}$, $n^{i}$] is denoted as $\bm{f}^{i}_{P}$.
Typically, there are multiple surrounding-view cameras on the autonomous vehicle, resulting in multi-view images. A 3D point can be projected onto one or two views. If the PoI is projected to only one view, this view will be leveraged to perform image feature sampling. Otherwise, we follow [5] to randomly choose one view for image feature sampling. Given a 3D PoI $P^{i}$ in the LiDAR coordinate system, the projection onto the 2D image plane can be computed as follows:
$P_{\text{image}}^{i}=\mathcal{I}\cdot(\mathcal{R}\cdot P^{i}+\mathcal{T}),$ (3)
where $P_{\text{image}}^{i}$ represents the coordinates of the projected PoI on the image plane, $\mathcal{I}$ is the intrinsic camera matrix, $\mathcal{R}$ is the rotation matrix of the extrinsic parameters, and $\mathcal{T}$ is the translation vector of the extrinsic parameters.
In our camera encoder, we exploit an image backbone network followed by an FPN [35] to produce feature maps at P2, P3, P4, and P5 (1/4, 1/8, 1/16, and 1/32 downsampling, respectively). For a given projected point $P_{\text{image}}^{i}$, a set of image feature $\{\bm{g}^{i}_{j}\}_{j=2}^{5}$ is sampled from the multi-scale feature maps via bilinear interpolation. The sampled multi-scale image features are then aggregated as:
$\bm{f}_{I}^{i}=\frac{\sum_{j=2}^{5}\text{exp}(w_{j}^{i})\cdot\bm{g}_{j}^{i}}{\sum_{j=2}^{5}\text{exp}(w_{j}^{i})},$ (4)
where $\bm{f}_{I}^{i}$ represents the aggregated image feature and $\{w^{i}_{j}\}_{j=2}^{5}$ are scale weight coefficients, which are predicted by a linear layer according to the query feature $\bm{v}$.
Figure 3: An illustration of our dynamic fusion block.
3.4 Dynamic Multi-Modal Feature Fusion
After feature sampling, each PoI is attached with a multi-modal feature pair. The next step is to fuse the multi-modal feature and integrate it into the object query. One simple solution is to exploit a static linear layer over the concatenated feature pair for fusion. However, using static linear layers ignores the fact that the image and point cloud features contribute differently to different objects. To this end, our fusion block capitalizes on a dynamic fusion scheme.
The detail of our dynamic fusion block is depicted in Figure 3. After PoI generation and feature sampling, each object query $\bm{Q}$ is represented by a set of PoIs $\{P^{i}\}$, and each PoI is attached with a sampled multi-modal feature pair $\{\bm{f}_{P}^{i},\bm{f}_{I}^{i}\}$. Firstly, we exploit conventional linear layers to produce dynamic fusion parameters $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$ according to the query feature $\bm{v}$. After that, the concatenated multi-modal feature pair of each PoI is individually integrated through the dynamic linear layers parameterized by $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$. Note that the distinction between the dynamic linear layer and the conventional linear layer is that the parameters of the dynamic one are produced on the fly. Subsequently, the features of PoIs derived from the same object query are concatenated together and fed into an additional linear layer to perform PoI feature aggregation. Finally, the aggregated feature is added back to the query feature. The dynamic linear layers and the conventional linear layer leveraged for PoI feature aggregation are followed by a Layer Normalization operation [1] and a ReLU activation layer.
3.5 Prediction Head and Training Objective
Prediction Head.
Our prediction head consists of a classification head and a sibling box regression head. In the classification head, we predict the binary classification score for each category. In the regression head, we follow the iterative refinement scheme [55, 84, 37] to predict the center delta over the predicted box center from the previous iteration. The other parts of the predicted 3D bounding box, including the box dimension, the heading direction, and the velocity, are independently predicted at each time.
Training Objective.
We follow the practice of set prediction [6] for target assignment. Specifically, we exploit bipartite matching to perform one-to-one assignments between the predicted 3D bounding boxes and the ground-truth boxes. In corresponding to the prediction head, our training objective also includes two parts: a focal loss [36] for the classification head and an L1 loss for the regression head. The overall training objective is computed as:
$L=\alpha\cdot L_{cls}(\bm{y}_{gt},\bm{y}_{pred})+\beta\cdot L_{reg}(\bm{b}_{gt},\bm{b}_{pred})$ (5)
where $\bm{y}_{pred}$ is the classification score, $\bm{b}_{pred}$ is the predicted 3D box, $\alpha$ and $\beta$ are set as 2.0 and 0.25 following [69, 40] to balance the losses. The training objective is applied to the predicted output of each recurrent iteration.
4 Experiments
Method Present at Modality Camera Backbone LiDAR Backbone val test
NDS $\uparrow$ mAP $\uparrow$ NDS $\uparrow$ mAP $\uparrow$
DETR3D [63] CORL’21 C ResNet-50 - 43.4 34.9 47.9 41.2
BEVFormer [33] ECCV’22 C ResNet-50 - 51.7 41.6 56.9 48.1
CenterPoint [76] CVPR’21 L - VoxelNet 66.8 59.6 67.3 60.3
TransFusion-L [2] CVPR’22 L - VoxelNet 70.1 65.1 70.2 65.5
${}^{\star}$  [77] NeurIPS’21 C+L ResNet-50 VoxelNet 70.8 67.1 70.5 66.4
PointAugmenting [58] CVPR’21 C+L ResNet-50 VoxelNet - - 71.0 66.8
FUTR3D [8] CVPRW’22 C+L ResNet-101 VoxelNet 68.3 64.5 - -
TransFusion [2] CVPR’22 C+L ResNet-50 VoxelNet 71.3 67.5 71.6 68.9
UVTR [29] NeurIPS’22 C+L ResNet-101 VoxelNet 70.2 65.4 71.1 67.1
BEVFusion (PKU) [34] NeurIPS’22 C+L Dual-Swin-T VoxelNet 72.1 69.6 73.3 71.3
DeepInteraction [75] NeurIPS’22 C+L ResNet-50 VoxelNet 72.6 69.9 73.4 70.8
BEVFusion (MIT) [42] ICRA’23 C+L Swin-T VoxelNet 71.4 68.5 72.9 70.2
${}^{\star}$ [26] CVPR’23 C+L ResNet-50 VoxelNet 72.0 69.3 74.0 71.5
${}^{\star}$ [31] arXiv’23 C+L ResNet-50 Sparse-ResUNet 72.7 70.4 74.0 70.6
UniTR [60] ICCV’23 C+L Multi-Modal DSVT Multi-Modal DSVT 73.1 70.0 74.1 70.5
SparseFusion [67] ICCV’23 C+L ResNet-50 VoxelNet 72.8 70.4 73.8 72.0
${}^{\triangle}$ [69] ICCV’23 C+L VoVNet-99 VoxelNet 72.9 70.3 74.1 72.0
ObjectFusion [5] ICCV’23 C+L Swin-T VoxelNet 72.3 69.8 73.3 71.0
PoIFusion (ours) - C+L ResNet-50 VoxelNet 73.2 71.2 74.3 72.6
PoIFusion (ours) - C+L Swin-T VoxelNet 73.6 71.7 74.9 73.4
Table 1: Comparison of 3D object detection performance on nuScenes dataset. “C” means camera data (RGB images), “L” means LiDAR data (point clouds), $\star$” indicates the image input depends on predicted 2D instance masks, and $\triangle$” indicates the image resolution is enlarged to $1600\times 640$. We highlight the top-2 entries with bold font. We do not apply any test-time augmentation or model ensemble.
4.1 Experimental Setup
Dataset and Metric.
We evaluate our method on the nuScenes dataset [4], which includes 700, 150, and 150 driving scenes for training, validation, and testing. The vehicle for data collection is equipped with a 32-beam LiDAR sensor and 6 surrounding-view RGB cameras, thus providing both point clouds and multi-view images. Besides, this dataset annotates more than 1.4 million 3D bounding boxes, across 10 common categories on the street. Both the multi-sensor data and annotation facilitate the exploration of multi-modal 3D object detection. We follow the official policy to mainly evaluate our method in terms of mean average precision (mAP) and nuScenes detection score (NDS). The mAP is averaged over distance thresholds $0.5m$, $1m$, $2m$, and $4m$ on the BEV across $10$ classes. NDS is a weighted average of mAP and other true-positive metrics including mATE, mASE, mAOE, mAVE, and mAAE.
Network Configuration.
The camera encoder is composed of an image backbone network (i.e., ResNet [23] or Swin-Transformer [41]) and an FPN [35] to enrich the multi-scale information. The LiDAR encoder consists of a 3D sparse VoxelNet [80, 70] and a BEV backbone network. We follow the common practice of setting the image resolution as $800\times 448$ and setting the voxel size as ($0.075m$, $0.075m$, $0.2m$). In our multi-modal object decoder, there are $900$ object queries. The feature dimension of query embedding is set as $256$. The channel of image and point cloud features is also transformed to $256$ before being fed into the multi-modal decoder. We equally divide the channels of feature maps of each modality into 4 groups, and generate PoIs for each group. This grouping operation improves the capacity of the network [66] and reduces the computation costs of the dynamic fusion block. Moreover, the first dynamic fusion layer halves the channel of the query embedding, and the second dynamic fusion layer restores the feature dimension. The multi-modal object decoder is iteratively applied 6 times, with shared parameters.
Training and Inference.
We follow the same training pipeline of previous methods [42, 5, 67, 34, 75], and use the same pretrained models. Specifically, the camera encoder is initialized with the model weights of a Mask R-CNN [24] model pre-trained on nuImage [4] dataset. The LiDAR encoder is initialized with the model weights of pretrained TransFusion-L [2]. Our fusion framework is trained for 6 epochs with the AdamW optimizer. The training sample is resampled by CBGS [82] strategy. The initial learning rate is set as 1e-4, adapted with the one-cycle learning rate policy, and the weight decay is set as 0.01. To speed up training convergence, the query denoising technique [39, 69, 38] is utilized. Data augmentation including random flip, random rotation, random translation, random scaling, and random modal masking is adopted. We exploit 8 GPUs for training, with 2 samples on each GPU.
At inference, PoIFusion outputs the top 300 detection boxes without NMS. We don’t exploit any test time augmentation or model ensemble techniques.
4.2 Comparison with Other Methods
Main Results of Detection.
In Table 1, we conduct a comprehensive comparison on the nuScenes 3D object detection benchmark. The methods are divided into three groups: camera-based (“C”), LiDAR-based (“L”), and multi-modal (“C+L”). Broadly speaking, the multi-modal methods outperform the methods that leverage only one single modality, validating the benefits of integrating the semantic-intensive RGB image and location-aware LiDAR point clouds.
Method
Modality
AMOTA (%) $\uparrow$
AMOTP (%) $\downarrow$
CenterPoint [76]
L
63.7
60.6
VoxelNeXt [11]
L
70.2
64.0
TransFusion [2]
C+L
71.8
60.3
BEVFusion (MIT) [42]
C+L
72.8
59.4
ObjectFusion [5]
C+L
74.2
54.3
PoIFusion(ours)
C+L
75.1 50.7
Table 2: Comparison of 3D multi-object tracking performance. The results are evaluated on the nuScenes validation set.
Method mAP NDS Latency (ms)
A100
V100
Transfusion [2]
67.5
71.3
153.8
190.3
BEVFusion (MIT) [42]
68.5
71.4
113.9 135.5
DeepInteraction [75]
69.9
72.6
204.1
344.8
CMT [69]
70.3
72.9
210.8
387.1
${}^{\dagger}$ [69]
70.3
72.9
159.7
PoIFusion (ours)
71.2 73.2
115.2
163.7
Table 3: Runtime comparison.normal-$\dagger$” means this model is accelerated with the Flash-Attention Operator [15]. NVIDIA V100 GPU is not compatible with the Flash Attention package.
Figure 4: Qualitative results on the nuScenes validation set. One example from a sunny day and another example from a rainy night are presented. Our PoIFusion precisely detects 3D objects under varying weather and lighting. We use different colors for different categories.
NDS
mAP
center only
72.8
70.6
center + corner
73.2 71.2
NDS
mAP
baseline
72.4
70.1
+ B.T.
72.8
70.8
+ B.T. & P.S.
73.2 71.2
NDS
mAP
static fusion
71.8
69.3
dynamic fusion
73.2 71.2
NDS
mAP
camera only
40.6
29.4
LiDAR only
70.2
65.3
camera + LiDAR
73.2 71.2
image resolution
NDS
mAP
800$\times$448
73.2
71.2
1600$\times$640
73.7 71.9
voxel size
NDS
mAP
(0.125, 0.125, 0.2)
71.6
69.1
(0.1, 0.1, 0.2)
72.3
69.8
(0.075, 0.075, 0.2)
73.2 71.2
Table 4: Ablative experiments. In these experiments, the backbone networks are ResNet-50 and VoxelNet. Unless specified, image resolution is set as $800\times 448$, and the voxel size of point clouds is ($0.075m$,$0.075m$,$0.2m$). Our default setting is highlighted in gray.
Our proposed PoIFusion has set a state-of-the-art record on the nuScenes dataset. Specifically, equipped with ResNet-50, our method achieves 73.2% NDS / 71.2% mAP on the validation set and achieves 74.3% NDS / 72.6% mAP on the test set. Upgrading the image backbone to a more powerful Swin-T [41], the performance on the test set boosts to 74.9% NDS and 73.4% mAP, which makes an absolute improvement of 0.8% NDS and 1.4% mAP over the previous best results [69]. Compared to the unified-view methods [34, 42, 29], our PoIFusion better keeps modal-specific information and thus outperforms the representative BEVFusion [42] by 2.0% NDS and 3.2% mAP on the nuScenes test set. In comparison with recent work, i.e., ObjectFusion [5], which also generates object-centric features, our method performs multi-modal fusion at adaptive PoIs, instead of fusing the regional features extracted from different representation spaces. The employment of PoIs improves the flexibility of sampling location and enables fine-grained feature fusion, thus working better. Moreover, the qualitative results are shown in Figure 4, with one example from a sunny day and another example from a rainy night. Our PoIFusion can robustly perform 3D object detection under varying weather and lighting conditions.
Extended Results of MOT.
In addition to the main results of 3D object detection, we evaluate the generalizability of our method on the 3D multiple object tracking task of the nuScenes dataset. Specifically, we follow CenterPoint [76] to adopt the “tracking-by-detection” scheme that offline links the detection boxes into tracking tubelets. Note that other compared methods, except VoxelNeXt [11], also follow this same scheme. As shown in Table 2, our PoIFusion works the best (75.1% AMOTA and 50.7%AMOTP) among the compared methods, demonstrating our better detection results can also benefit other downstream tasks.
Runtime Comparison.
Furthermore, we conduct runtime comparisons among the recent works. For a fair comparison, we re-evaluate the latency on an NVIDIA A100 and an NVIDIA V100 GPU. Since the implementation of the voxelization operation is a confounder for runtime evaluation but will not influence the detection result, we exclude its time cost in comparison. As shown in Table 3, the latency of our PoIFusion is comparable to the fastest method, i.e., BEVFusion (MIT), which further shows the potential of our method to serve as a strong baseline for future investigation.
4.3 Ablative Experiments
Effect of Components in the Multi-Modal Decoder.
In this section, we analyze the design of the proposed multi-modal decoder from three aspects: (a) anchor points, (b) operations to generate PoIs, and (c) the fusion block. As shown in Table LABEL:tab:ablation:anchor, the model that only uses the center point achieves 72.8% NDS and 70.6% mAP. By additionally involving the corner points, the NDS and mAP boosts to 73.2% and 71.2%, respectively. This comparison shows the benefits of representing the geometric property of query boxes. In Table LABEL:tab:ablation:transformation, we validate the operations to derive PoIs from anchor points. The baseline directly takes the anchor points as PoIs without adaptation, achieving 70.1% mAP. By involving the holistic box-level transformation and further applying the point-wise shift, the performance boosts to 70.8 % mAP and 71.2% mAP, respectively. The online adjustment of PoIs ameliorates the misalignment in feature sampling and thereby boosts the detection result.
Table LABEL:tab:ablation:fusion_block compares the effect of different fusion schemes, i.e., static fusion versus dynamic fusion. In static fusion, two conventional linear layers are exploited to process the concatenated multi-modal features at each PoI. The dynamic fusion scheme refers to our proposed on in Section 3.4. Thanks to the adjustment of fusion layers’ parameters for each individual object query, our dynamic fusion outperforms the static setting by 1.4% NDS and 1.9% mAP.
Effect of Different Modalities.
In this experiment, we train models with single-modality input following the same setting as our default one. As shown in Table LABEL:tab:ablation:modality, the image-only baseline achieves 40.6% NDS and 29.4% mAP. Meanwhile, the LiDAR-only baseline achieves 70.2% NDS and 65.3% mAP, which is similar to TransFusion-L [2] (70.1% NDS and 65.1% mAP). Integrating the multi-modal information remarkably boosts the performance to $73.2\%$ NDS and $71.2\%$ mAP, validating the capability of our proposed fusion paradigm to integrate the complementary information from images and point clouds.
Effect of the Input Resolution.
The input resolution is also one of the key factors that influence the detection result. In Table LABEL:tab:ablation:img_resolution, we enlarge the image resolution from $800\times 448$ to $1600\times 640$, obtaining 0.5% NDS and 0.7% mAP gain. In Table LABEL:tab:ablation:voxel_size, we further analyze the influence of the voxel size. Either a larger image resolution or a smaller voxel size facilitates the feature representation of the corresponding modality, thus boosting the detection result. Our default setting employs an image resolution of $800\times 448$ and a voxel size of ($0.0075m,0.0075m,0.2m$), which is the same as most of recent works [2, 5, 67, 75].
Offsets (m)
0.0
0.2
0.4
0.6
0.8
1.0
NDS (%)
73.2
73.0
72.9
72.7
72.3
72.0
mAP (%)
71.2
71.1
70.8
70.5
70.0
69.4
Table 5: Sensor misalignment. Performance comparison on the impact of translation offsets due to calibration errors.
4.4 Detection Robustness Analysis
Sensor Misalignment.
In real applications, sensors can be misaligned due to physical impacts, installation errors, time-related drift, or some other unexpected reasons. This experiment investigates the robustness of PoIFusion against sensor misalignment. Specifically, we follow [2] to randomly add translation offsets to the calibration matrix at inference. The errors are formulated as uniformly distributed noise, with the maximum offset value varying from $0.2m$ to $1.0m$. As shown in Table 5, even when the maximum translation offset is $1.0m$, our method can achieve 72.0% NDS and 69.4% mAP, demonstrating the robustness of our fusion paradigm against sensor misalignment.
Sensor Failure.
In addition to sensor misalignment, we further explore a more serious condition, i.e., sensor failure. To validate our method under camera failure cases, we randomly drop several images by filling the corresponding images as all zeros. The results with different numbers of dropped images are presented in Table 6. The performance keeps dropping along with the increasing number of cameras that fail to work. Even with half of the cameras blocked, our PoIFusion achieves 68.6% mAP, obviously better than the LiDAR-only method. Moreover, in the extreme case that all of the cameras are not working, our method works comparable to the LiDAR-only baseline.
# Dropped Images
0
1
3
6
NDS (%)
73.2
72.8
72.0
70.2
mAP (%)
71.2
70.3
68.6
65.6
Table 6: Camera failure. Performance comparison on the impact of dropped images due to camera failure.
Discarded Sectors
${}^{\circ}$
${}^{\circ}$
${}^{\circ}$
${}^{\circ}$
${}^{\circ}$
NDS (%)
73.2
72.7
72.0
71.3
70.7
mAP (%)
71.2
70.2
68.9
67.6
66.4
Table 7: LiDAR failure. Performance comparison on the impact of discarded point clouds within a sector due to LiDAR failure.
Furthermore, to simulate the failure of a LiDAR sensor, we randomly discard point cloud data within a specified sectorial region based on the azimuthal angle. As shown in Table 7, in general, the impact of LiDAR failure is more pronounced than that of camera failure, since the spatial information captured by the LiDAR sensor is essential for precise localization. Notwithstanding the total absence of point cloud data within a $24^{\circ}$ sector, the mAP achieved by our method is still 1.1% higher than the LiDAR-only baseline. These two experiments demonstrate the robustness of our proposed PoIFusion against Sensor Failure.
5 Conclusion
In this work, we present PoIFusion, a query-based multi-modal 3D object detector, which is characterized by performing multi-modal feature sampling and fusion at generated points of interest. On the one hand, PoIFusion keeps the original view of the image and point cloud feature to preserve the modal-specific information. On the other hand, the employment of points of interest facilitates fine-grained fusion and enables representing object features in a flexible way. Our PoIFusion sets a state-of-the-art record on the highly competitive nuScenes dataset, and in the meantime keeps efficient runtime. Moreover, the experimental analysis demonstrates our method is robust to sensor misalignment and failure. We suppose our PoIFusion would serve as a strong baseline with huge potential for future exploration.
Supplementary Material
This supplementary material is composed of three sections. In Section A, we introduce supplementary details of our implementation. Then, in Section B, we provide additional experiments to further validate the merits of our proposed method. Finally, in section C, we visualize more examples to help intuitively present the detection results.
Appendix A Supplementary Details
Query Initialization.
As introduced in our main paper, we formulate each object query as a learnable dynamic 3D box $\bm{b}\in\mathbb{R}^{8}$ and an attached feature vector $\bm{v}\in\mathbb{R}^{256}$. Specifically, the query box is represented as:
$\bm{b}=[x_{c},y_{c},z_{c},w,l,h,sin\theta,cos\theta],$ (6)
where [$x_{c}$, $y_{c}$, $z_{c}$] is the center location, [$w$, $l$, $h$] is the box dimension, and $\theta$ is the azimuth angle. Notably, while velocity is a component of the detection results, it is not factored into the query boxes. When we initialize the query box, the BEV center location [$x_{c}$, $y_{c}$] is set to be uniformly distributed on the BEV, and the height center $z_{c}$ is set as 0. Besides, the dimensions $w$, $l$, $h$ of each box are respectively initialized as 6, 3, 2, and the azimuth is initialized as 0.
Distance-Biased Self-Attention Layer.
As outlined in Section 3.1, the multi-modal decoder utilizes a distance-biased self-attention layer [38] to capture the relationships and dependencies between different object queries. Here, we provide a comprehensive review of it.
Specifically, a conventional self-attention layer [56] calculates the attention score based on the dot product of query and key, which is detailed as:
[…]
where Q and K correspond to the query and key, and C is the channel dimension. The distance-biased self-attention layer incorporates a bias term that is contingent upon the centroid distance between pairs of query boxes. The adjusted attention score is computed as:
[…]
where $D$ indicates the centroid distance between query boxes, and $\psi$ is a scalar produced by a linear layer according to the query feature. In our multi-modal object decoder, we use 8 heads in the self-attention layer. The $\psi$ is not shared in different heads. We emphasize that the distance-biased self-attention layer is not the contribution of our work. We implement the self-attention layer with it since it improves the adaptability of the model with negligible costs.
Appendix B Additional Experimental Analysis
Different Weathers and Lighting Conditions.
The 3D object detection system on autonomous driving vehicles is supposed to be capable of working under varying lighting and weather conditions. In the main manuscript, we have shown two qualitative examples to give a glimpse of the robustness of our PoIFusion. In this experiment, we further follow [42] to divide the validation set of the nuScenes dataset into two pairs of subsets: sunny/rainy and day/night for comprehensive quantitative evaluation. The evaluated performance in terms of mAP (%) is summarized in Table S1. In general, the multi-modal 3D object detectors are not likely to be affected by the rainy weather. The changing of lighting condition significantly affects the detection performance, i.e., the mAP on the night subset is much lower than that of the day subset. Among the compared methods, our PoIFusion achieves the best performance under different weather and lighting conditions, further demonstrating the robustness of our fusion strategy.
mAP (%) Modality Sunny Rainy Day Night
BEVDet C 32.9 33.7 33.7 13.5
Centerpoint L 62.9 59.2 62.8 35.4
MVP C+L 65.9 66.3 66.3 38.4
BEVFusion C+L 68.2 69.9 68.5 42.8
PoIFusion (ours) C+L 71.1 70.2 71.3 45.0
Table S1: Different lighting and weather conditions. The reported performance is compared in terms of mAP (%) on the nuScenes validation set. We report the result of our PoIFusion with the Swin-T image backbone network.
Effect of Iteration Times of the Multi-Modal Decoder.
As introduced in Section 3.1, our multi-modal object decoder is iteratively applied to perform feature fusion and box refinement. In Table S2, we analyze how the iteration times of our multi-modal decoder influence the 3D object detection performance. Specifically, we vary the iteration times from 1 to 8 times and re-train the model for evaluation. In the beginning, the performance consistently improves along with the increasing iteration time. The peak performance is achieved when we apply the multi-modal object decoder 6 times. When we further increase the iteration time to 8, the performance drops for 0.5% NDS and 0.9% mAP. Therefore, in our default configuration, the multi-modal object decoder is iteratively applied 6 times. Moreover, as shown in the table, our method provides the potential to make a good trade-off between accuracy and latency. Specifically, when the iteration time is reduced to 4, our PoIFusion can still achieve the highest NDS (73.1%) and mAP (70.7%) among the compared methods presented in Table 1, and in the meantime run faster than the fastest competitor (i.e., BEVFusion) presented in Table 3.
# Iteration Times
NDS (%)
mAP (%)
Latency (ms)
1
67.1
63.4
87.7
2
71.9
69.3
94.2
4
73.1
70.7
99.8
6
73.2 71.2
115.2
8
72.7
70.3
119.2
Table S2: Effect of iteration times of the multi-modal decoder. The reported performance is compared on the nuScenes validation set. We exploit ResNet-50 as the image backbone network. The latency is evaluated on an NVIDIA A100 GPU.
Shared Parameters VS. Unshared Parameters
of the Multi-Modal Decoder. Since the multi-modal object decoder is iteratively applied multiple times, we can exploit either shared parameters or unshared parameters. The results of using shared parameters and unshared parameters are presented in Table S3. As shown in this comparison, these two models achieve comparable performance. Moreover, sharing parameters of the multi-modal object decoder makes a smaller model size. Therefore, we use shared parameters of the multi-modal decoder by default.
NDS (%)
mAP (%)
Shared Parameters
73.2
71.2
Unshared Parameters
73.3
70.9
Table S3: Shared versus unshared parameters of the multi-modal object decoder. The models are evaluated on the nuScenes validation set, with ResNet-50 as the image backbone network.
Training without the pretrained LiDAR encoder.
In our main experiment, we follow the training policy of most recent works [2, 42, 67, 5, 75], i.e., starting with the model weight of pretrained TransFusion-L and training the fusion model for 6 epochs. Here, we further study whether our PoIFusion can also be trained without a pretrained TransFusion-L model. Specifically, in this experiment, we follow the training policy of CMT [69] that trains the fusion framework with 20 epochs, with only the pretrained image backbone network. The performance comparison between our PoIFusion and CMT is presented in Table S4. Please note that these two models are trained with the same training policy, backbone network, and input resolution (details are summarized in the caption of this table). As shown in the table, our PoIFusion achieves 72.3% NDS and 68.4% mAP without pretrained LiDAR encoder, outperforming CMT with the same input resolution and training policy by 1.5% NDS and 0.5% mAP. This experiment demonstrates our PoIFusion works very well under various training policies, further exhibiting its potential to serve as a strong baseline for future research.
Method
NDS (%)
mAP (%)
CMT
70.8
67.9
PoIFusion (ours)
72.3 68.4
Table S4: Training without a pretrained TransFusion-L model. The compared models are trained with the same training policy and evaluated on the nuScenes validation set. Both models use ResNet-50 and VoxelNet as the backbone network, with image resolution set as $800\times 320$ and voxel size set as $(0.1m,0.1m,0.2m)$.
Appendix C Additional Qualitative Results
In our main manuscript, we only visualize two scenes due to the limitation of pages. Here, we add more qualitative examples in Figure S1 to help intuitively present the detection results of our method. From the visualized results, it can be seen that our proposed method is able to accurately detect 3D objects under complex road conditions.
Figure S1: Supplement of Qualitative results on the nuScenes validation set.

### v2

@@
-PoIFusion: Multi-Modal 3D Object Detection via Fusion at Points of Interest
+Feras Dayoub
 Abstract
-In this work, we present PoIFusion, a simple yet effective multi-modal 3D object detection framework to fuse the information of RGB images and LiDAR point clouds at the point of interest (abbreviated as PoI).
-Technically, our PoIFusion follows the paradigm of query-based object detection, formulating object queries as dynamic 3D boxes.
-The PoIs are adaptively generated from each query box on the fly, serving as the keypoints to represent a 3D object and play the role of basic units in multi-modal fusion.
+In this work, we present PoIFusion, a conceptually simple yet effective multi-modal 3D object detection framework to fuse the information of RGB images and LiDAR point clouds at the points of interest (PoIs).
+Different from the most accurate methods to date that transform multi-sensor data into a unified view or leverage the global attention mechanism to facilitate fusion, our approach maintains the view of each modality and obtains multi-modal features by computation-friendly projection and interpolation.
+In particular, our PoIFusion follows the paradigm of query-based object detection, formulating object queries as dynamic 3D boxes and generating a set of PoIs based on each query box.
+The PoIs serve as the keypoints to represent a 3D object and play the role of the basic units in multi-modal fusion.
 Specifically, we project PoIs into the view of each modality to sample the corresponding feature and integrate the multi-modal features at each PoI through a dynamic fusion block.
@@
 Our approach prevents information loss caused by view transformation and eliminates the computation-intensive global attention, making the multi-modal 3D object detector more applicable.
-We conducted extensive experiments on the nuScenes dataset to evaluate our approach.
-Remarkably, our PoIFusion achieves 74.9% NDS and 73.4% mAP, setting a state-of-the-art record on the multi-modal 3D object detection benchmark.
-Codes will be made available via https://djiajunustc.github.io/projects/poifusion.
+We conducted extensive experiments on nuScenes and Argoverse2 datasets to evaluate our approach.
+Remarkably, the proposed approach achieves state-of-the-art results on both datasets without any bells and whistles, i.e., 74.9% NDS and 73.4% mAP on nuScenes, and 31.6% CDS and 40.6% mAP on Argoverse2.
+The code will be made available at https://djiajunustc.github.io/projects/poifusion.
+Keywords:
+3D Object Detection Autonomous Driving Multi-Sensor Fusion
 1 Introduction
+Figure 1: A comparison between the representative multi-modal fusion mechanism in the literature and ours: (a) fusion with a unified view, (b) fusion with global attention, and (c) our proposed fusion at points of interest.
+“View Trans.”: view transformation.
+“PoI”: points of interest.
 Autonomous vehicles are usually equipped with an array of sensors to facilitate safe driving, among which cameras and LiDARs are the most popular.
 These two sensors are complementary to each other: cameras provide rich textual and color information, while LiDAR sensors supply precise spatial measurements.
-The effective camera-LiDAR data fusion is widely regarded as a promising direction to achieve high-quality 3D object detection [2, 34, 42, 64], which has attracted a surge of research interest in the community.
-Figure 1: A comparison of (a) fusion with unified view, (b) fusion with global attention, and (c) our PoIFusion. “VT” and “PC” indicate view transformation and point clouds, respectively.
+The effective camera-LiDAR data fusion is widely regarded as a promising direction to achieve high-quality 3D object detection Bai et al.
+(2022); Liang et al.
+(2022); Liu et al.
+(2023b); Wang et al.
+(2023d); Zhang et al.
+(2024); Mao et al.
+(2023), which has attracted a surge of research interest in the community.
 The fundamental challenge of camera-LiDAR data fusion arises from the discrepancy of their representation space (i.e., 2D perspective view versus 3D space).
-To ameliorate this challenge, one common solution is to transform the image and point cloud representations into a unified bird-eye view [42, 34, 22], as depicted in Figure 1(a), or into 3D space [29].
-Another recent approach [2, 69], as shown in Figure 1(b), keeps the representation in its original view and abstracts multi-modal features into object queries [6, 40] with the global attention mechanism [56].
+To ameliorate this challenge, one common solution is to transform the image and point cloud representations into a unified bird-eye view Liu et al.
+(2023b); Liang et al.
+(2022); Ge et al.
+(2023), as depicted in Figure 1(a), or into 3D space Li et al.
+(2022a).
+Another recent approach Bai et al.
+(2022); Yan et al.
+(2023), as shown in Figure 1(b), keeps the representation in its original view and abstracts multi-modal features into object queries Carion et al.
+(2020); Liu et al.
+(2022) with the global attention mechanism Vaswani et al.
+(2017).
 However, both of these approaches have inherent issues.
@@
 However, depth estimation is an error-prone task, with errors having a deleterious effect on any downstream tasks (such as object recognition).
-Moreover, the direct grid-to-grid fusion [34, 42] loses a significant portion of the original representational strengths which comprises modal-specific information [75].
+Moreover, the direct grid-to-grid fusion Liang et al.
+(2022); Liu et al.
+(2023b) loses a significant portion of the original representational strengths which comprises modal-specific information Yang et al.
+(2022c).
 The second query-based approach can avoid feature ambiguity and information loss by keeping the original view of the feature representation.
 However, the adoption of global attention to integrating multi-modal features incurs high computation and memory overhead.
-For instance, the state-of-the-art algorithm [69], which follows the query-based approach of fusion with global attention, relies on the well-optimized Flash Attention operator [15] to cut down the time and memory consumption.
+For instance, the state-of-the-art algorithm Yan et al.
+(2023), which follows the query-based approach of fusion with global attention, relies on the well-optimized Flash Attention operator Dao et al.
+(2022) to cut down the time and memory consumption.
 The high overhead has become an obstacle that hinders the wide application of the algorithm.
-Furthermore, it is difficult to extract the object-relevant feature with the global attention mechanism [84, 21], especially when it comes to such a large 3D space like the autonomous driving scenario.
+Furthermore, it is difficult to extract the object-relevant feature with the global attention mechanism Zhu et al.
+(2020); Gao et al.
+(2022), especially when it comes to such a large 3D space like the autonomous driving scenario.
 The benefit of the query-based approach inspires us to preserve the original view of features from each modality, while the above two issues motivate our exploration of an alternative paradigm to replace dense feature interaction of the global attention with sparse point projection and feature sampling, as illustrated in Figure 1(c).
@@
 In this manuscript, these representative points are referred to as Points of Interest (PoIs), and our framework is named PoIFusion.
-Intuitively, a naive way to represent a 3D query box with points is to use the center point [8, 69, 40].
+Intuitively, a naive way to represent a 3D query box with points is to use the center point Chen et al.
+(2022a); Yan et al.
+(2023); Liu et al.
+(2022).
 However, simply representing a 3D box with its center point totally ignores the geometric properties, such as the size and rotation angle.
-For supplementary, an improved design is to involve the corners [49].
+For supplementary, an improved design is to involve the corners Sheng et al.
+(2021).
 Nevertheless, sampling multi-modal features according to the projected location of center and corner points also incurs the problem of feature misalignment, since the projection of a 3D box may not be a tight bounding box onto the image view.
-The issue remains even if a box is well-located in the 3D space [5], not to mention that the query box is not guaranteed to be accurately localized.
+The issue remains even if a box is well-located in the 3D space Cai et al.
+(2023), not to mention that the query box is not guaranteed to be accurately localized.
 To ameliorate this problem, our PoIs are adaptively generated from the center and corner points, with box-level and point-level transformation parameters online predicted according to the query feature.
@@
 The proposed adaptive PoIs, together with the dynamic fusion block, enable our PoIFusion to efficiently sample the object-relevant multi-modal features and make the best of modal-specific information, thus improving 3D object detection.
-We benchmark the proposed PoIFusion framework on the nuScene dataset and conduct extensive experiments to validate our designs.
-Notably, our method sets a new record on this challenging dataset, achieving 74.9% NDS and 73.4% mAP on the test set without bells and whistles.
-In summary, our contributions are three-fold:
+We evaluate the proposed PoIFusion framework on the nuScenes and Argoverse2 datasets, conducting a comprehensive set of experimental analyses to validate our design choices.
+Notably, PoIFusion achieves state-of-the-art performance on the highly competitive nuScenes benchmark, attaining 74.9% NDS and 73.4% mAP on the test set without any bells and whistles.
+Moreover, when applied to the more challenging Argoverse2 dataset, our method achieves 40.6% mAP and 31.6% CDS, improving the best performance in the literature by absolutely 4.5% mAP and 3.8% CDS.
+In summary, we make three-fold our contributions:
   • We propose a novel PoIFusion framework for multi-modal 3D object detection that preserves modality-specific representation spaces while efficiently extracting and fusing features through sparse interactions.
   • We present the design of fusion at points of interest, conveying an elegant view that the entity involved in the fusion module can be very flexible.
-  • We conduct extensive experiments to validate the effectiveness of our method, demonstrating its potential to serve as a strong baseline for future investigation.
+  • We conduct extensive experiments to validate the effectiveness of our method, demonstrating its potential to serve as a strong baseline for this field.
 2 Related Work
-LiDAR-Based 3D Object Detection.
+2.1 LiDAR-Based 3D Object Detection.
 LiDAR sensors capture 3D point clouds, which provide accurate spatial information that can be important for 3D object detection.
 Broadly, 3D object detection algorithms operated on point clouds can be categorized into two groups: point-based and voxel-based ones.
-Point-based 3D object detection algorithms [51, 74, 73] make direct use of the precise coordinates of points, progressively sampling keypoints and extracting local information with set abstraction operators [46, 47, 52, 53] that aggregate information form a point and its neighbours.
-In contrast, voxel-based 3D object detection algorithms [80, 72, 28, 16, 76, 50, 20, 17] first “voxelize” the point cloud by binning points into a regulare grid.
-The voxel representation enables straightforward feature extraction using standard (or sparse) convolutions [70, 9, 10] or sparse voxel Transformers [19, 59, 81, 27, 12, 18, 44].
-Figure 2: An overview of our proposed PoIFusion framework, which is mainly composed of (a) an image feature module, (b) a point cloud feature module, and (c) a multi-modal object decoder.
-In our PoIFusion, we independently extract the feature of each modality and keep the original representation view (i.e., image feature in the perspective view, and point cloud feature in the bird-eye view).
-The multi-modal feature maps are taken as the input of our multi-modal object decoder.
-The object decoder is iteratively applied 6 times to integrate the multi-modal feature sampled by generated points of interest (PoIs) and refine the object queries.
-Camera-Based 3D Object Detection.
+Point-based 3D object detection algorithms Shi et al.
+(2019); Yang et al.
+(2020, 2019) make direct use of the precise coordinates of points, progressively sampling keypoints and extracting local information with set abstraction operators Qi et al.
+(2017a, b); Shi et al.
+(2020a, 2023) that aggregate information form a point and its neighbors.
+In contrast, voxel-based 3D object detection algorithms Zhou and Tuzel (2018); Yang et al.
+(2022a); Lang et al.
+(2019); Deng et al.
+(2021a); Yin et al.
+(2021a); Shi et al.
+(2022); Fan et al.
+(2022b); Deng et al.
+(2021b); Yang et al.
+(2022b); Wang et al.
+(2023c, 2024) first “voxelize” the point cloud by binning points into a regular grid.
+The voxel representation enables straightforward feature extraction using standard (or sparse) convolutions Yan et al.
+(2018); Chen et al.
+(2022b, c); Shi et al.
+(2020b) or sparse voxel Transformers Fan et al.
+(2022a); Wang et al.
+(2023a); Zhou et al.
+(2022); Lai et al.
+(2023); Chen et al.
+(2023b); Dong et al.
+(2022); Mao et al.
+(2021).
+Figure 2: An overview of our proposed PoIFusion framework, which is mainly composed of (a) a camera feature module, (b) a LiDAR feature module, and (c) a PoI multi-modal decoder.
+In our method, we first independently extract the feature of each modality and keep the original representation view (i.e., image feature in the perspective view, and point cloud feature in the bird-eye view).
+The multi-modal feature maps are taken as the input of our PoI multi-modal decoder.
+The decoder is iteratively applied 6 times to integrate the multi-modal feature sampled with generated points of interest (PoIs) and to refine the object queries.
+In this figure, “proj.” stands for “projection”.
+2.2 Camera-Based 3D Object Detection.
 3D object detection is one of the oldest Computer Vision problems and a variety of approaches have been proposed over many years (we do not survey these here).
 For the purposes of 3D object detection from a moving car a popular approach, and one that can be readily fused with point-cloud estimates, is to measure detection success in the Birds-eye View (BEV) space.
-Earlier attempts along these lines – e.g., geometric uncertainty [43] and pseudo-LiDAR representation [62] – mainly focus on monocular 3D object detection [61, 48, 3, 14, 79].
-Autonomous driving vehicles are typically equipped with multiple cameras, providing perception information over the full 360 degrees.
+Earlier attempts along these lines – e.g., geometric uncertainty Lu et al.
+(2021) and pseudo-LiDAR representation Wang et al.
+(2019) – mainly focus on monocular 3D object detection Wang et al.
+(2021b); Reading et al.
+(2021); Brazil and Liu (2019); Chong et al.
+(2022); Zhang et al.
+(2023); Zhou et al.
+(2021).
+Autonomous driving vehicles are typically equipped with multiple cameras Xie et al.
+(2021); Guo et al.
+(2024), providing perception information over the full 360 degrees.
 To leverage the relationship between multi-view images, BEV-based and query-based algorithms have been explored.
-BEV-based 3D object detection [25, 33, 32, 71] explicitly performs view transformation to unify multi-view images into bird-eye-view representation.
-Query-based 3D object detection [40, 63, 38, 65, 54, 68] follows the pipeline of DETR [6], capitalizing on learnable object queries to extract multi-view information without view transformation.
-Multi-Modal 3D Object Detection.
+BEV-based 3D object detection Huang et al.
+(2021); Li et al.
+(2022c, 2023b, 2023a) explicitly performs view transformation to unify multi-view images into bird-eye-view representation.
+Query-based 3D object detection Liu et al.
+(2022); Wang et al.
+(2022); Liu et al.
+(2023a); Wang et al.
+(2023e); Shu et al.
+(2023); Xiong et al.
+(2023) follows the pipeline of DETR Carion et al.
+(2020), capitalizing on object queries to extract multi-view information without view transformation.
+2.3 Multi-Modal 3D Object Detection.
 Although the exploration of each individual modality has made encouraging progress, the accuracy and robustness of detection algorithms are still insufficient for safe driving.
-The fact that images and point clouds are naturally complementary to each other (i.e., rich semantic information versus precise spatial information) has motivated further exploration in multi-modal 3D object detection [60, 2, 8, 22, 67, 31, 26].
-In the early stage, the multi-modal approaches utilize point clouds as the principal component, introducing image features at the point level [58, 57, 77, 13] or proposal level [83, 7, 78] to enhance the features of the point clouds.
-Recently, inspired by multi-view 3D object detection, a series of works [34, 42, 29] propose that unifying the representation space with explicit view transformation [45] facilitates multi-modal fusion.
-Another group of methods [30, 2, 69, 75] leverages the attention mechanism in Transformer architecture [56] to perform multi-modal fusion in a sequential [2] or a parallel [69] manner.
-A recent work, ObjectFusion [5], fuses multi-modal features in a two-stage pipeline.
-It first generates region proposals with the image-augmented BEV features [2], and then extracts the object-centric feature [24, 16] from the voxel, image, and BEV space for further fusion and refinement.
-In this work, the proposed PoIFusion is also object-centric.
-However, in contrast to integrating the corresponding region-level feature from multiple modalities, we utilize the PoIs as the basic units to perform multi-modal fusion, which is adaptively derived from the query box - this element stands as the cornerstone of our approach, fundamentally improving the efficacy and flexibility of our framework.
+The fact that images and point clouds are naturally complementary to each other (i.e., rich semantic information versus precise spatial information) has motivated further exploration in multi-modal 3D object detection Wang et al.
+(2023b); Bai et al.
+(2022); Chen et al.
+(2022a); Ge et al.
+(2023); Xie et al.
+(2023); Li et al.
+(2024); Jiao et al.
+(2023); Yin et al.
+(2024).
+In the early stage, the multi-modal approaches utilize point clouds as the principal component, introducing image features at the point level Wang et al.
+(2021a); Vora et al.
+(2020); Yin et al.
+(2021b); Chen et al.
+(2022d) or proposal level Zhu et al.
+(2022); Chen et al.
+(2017); Yoo et al.
+(2020) to enhance the features of the point clouds.
+Recently, inspired by multi-view 3D object detection, a series of works Liang et al.
+(2022); Liu et al.
+(2023b); Li et al.
+(2022a) propose that unifying the representation space with explicit view transformation Philion and Fidler (2020) facilitates multi-modal fusion.
+Another group of methods Li et al.
+(2022b); Bai et al.
+(2022); Yan et al.
+(2023); Yang et al.
+(2022c) leverages the attention mechanism in Transformer architecture Vaswani et al.
+(2017) to perform multi-modal fusion in a sequential Bai et al.
+(2022) or a parallel Yan et al.
+(2023) manner.
+A recent work, ObjectFusion Cai et al.
+(2023), fuses multi-modal features in a two-stage pipeline.
+It first generates region proposals with the image-augmented BEV features Bai et al.
+(2022), and then extracts the object-centric feature He et al.
+(2017); Deng et al.
+(2021a) from the voxel, image, and BEV space for further fusion and refinement.
+In this work, the proposed PoIFusion scheme is also object-centric.
+However, in contrast to integrating the corresponding region-wise feature from multiple modalities, we adaptively generate PoIs from each object query, and leverage the PoIs as the basic units to perform multi-modal fusion - this element stands as the cornerstone of our approach, fundamentally improving the efficacy and flexibility of our multi-modal 3D object detection framework.
 3 Our Approach
@@
 3.1 Overview
-As illustrated in Figure 2, our PoIFusion consists of three main components: (a) an image feature module, (b) a point cloud feature module, and (c) a multi-modal object decoder.
-Given multi-view images and point clouds, perspective-view image feature maps $\bm{F}_{I}$ and bird-eye-view (BEV) point cloud feature maps $\bm{F}_{P}$ are independently extracted with the camera encoder and the LiDAR encoder.
-The camera encoder is composed of an image backbone network and an FPN [35] to obtain multi-scale features.
-The LiDAR encoder exploits the VoxelNet [70, 80] and a BEV backbone network, following the common practice of the voxel-based paradigm.
-Then, the multi-modal object decoder, which works as the core component in our method, is iteratively applied 6 times, progressively integrating the multi-modal features and refining the detection boxes.
+As illustrated in Figure 2, our PoIFusion consists of three main components: (a) a camera feature module, (b) a LiDAR feature module, and (c) a PoI multi-modal decoder.
+Given multi-view images and point clouds, perspective-view image feature maps $\bm{F}_{I}$ and bird-eye-view (BEV) point cloud feature maps $\bm{F}_{P}$ are independently extracted with the image encoder and the voxel encoder.
+The image encoder is composed of an image backbone network and an FPN Lin et al.
+(2017a) to obtain multi-scale features.
+The voxel encoder exploits a sparse 3D backbone network Yan et al.
+(2018); Zhou and Tuzel (2018) and a BEV backbone network, following the common practice of the voxel-based paradigm.
+Then, the PoI multi-modal decoder, which works as the core component in our method, is iteratively applied 6 times, progressively integrating the multi-modal features and refining the detection boxes.
 In our multi-modal decoder, each object query $\bm{Q}$ is formulated as an adaptive 3D box $\bm{b}_{t}$, associating with a feature vector $\bm{v}_{t}$ ($t$ indicates the iteration time step).
 In each iteration, the object queries are first fed into a self-attention layer to capture the relationships and dependencies between different objects.
-Here, we follow [38] to exploit a distance-biased self-attention layer in our decoder.
+Here, we follow Liu et al.
+(2023a) to exploit a distance-biased self-attention layer in our decoder.
 Subsequently, for each object query, the box-wise transformation parameters and point-wise shift parameters are generated based on the query feature $\bm{v}_{t}$, and applied on the query box $\bm{b}_{t}$ to obtain a set of PoIs $\bm{P}_{t}=\{P^{i}\}$.
@@
 Thirdly, the extracted multi-modal features at each PoI are dynamically integrated using a dynamic fusion block, followed by a feedforward network (FFN).
-Finally, a classification head and a regression head are applied to obtain the detection boxes, and the query feature and query box are updated, respectively.
+Finally, a classification head and a regression head are applied for prediction, and the query feature and query box are updated, respectively.
+3.2 Object Query Initialization.
+We formulate each object query as a learnable dynamic 3D box $\bm{b}\in\mathbb{R}^{8}$ and an attached feature vector $\bm{v}\in\mathbb{R}^{256}$. Specifically, the query box is represented as:
+$\bm{b}=[x_{c},y_{c},z_{c},w,l,h,sin\theta,cos\theta],$ (1)
+where [$x_{c}$, $y_{c}$, $z_{c}$] is the center location, [$w$, $l$, $h$] is the box dimension, and $\theta$ is the azimuth angle.
+Notably, although velocity is also predicted in the detection results, it is not factored into the query boxes.
+Prior to training, the object query is initialized as follows: the BEV center location [$x_{c}$, $y_{c}$] is set to be uniformly distributed on the BEV space, and the height center $z_{c}$ is set as 0.
+Besides, the dimension triplet [$w$, $l$, $h$] of each box is initialized as [6, 3, 2], which is about the average size of objects calculated on the dataset.
+and the azimuth is initialized as 0.
+The corresponding attached feature vector is randomly initialized.
 3.3 PoI Generation
@@
 $x_{c}^{\prime}$ $=x_{c}+t_{x},$ $w^{\prime}$ $=w\cdot\exp(t_{w}),$ (2)
-$y_{c}^{\prime}$ $=y_{c}+t_{y},$ $l^{\prime}$ $=l\cdot\exp(t_{l}),$
-$z_{c}^{\prime}$ $=z_{c}+t_{z},$ $h^{\prime}$ $=h\cdot\exp(t_{h}),$
-$\sin\theta^{\prime}$ $=\sin\theta+t_{\sin},$ $\cos\theta^{\prime}$ $=\cos\theta+t_{\cos}.$
-Once the box transformation has been applied, the center point and 8 corner points of $\bm{b^{\prime}}$ are collected together as the anchor points $\{A^{i}\}_{i=0}^{8}$.
-Finally, the point-wise shift is independently conducted on each anchor point $A^{i}$ to produce a PoI $P^{i}$ ($P^{i}=A^{i}+\Delta_{P}^{i}$).
+$y_{c}^{\prime}$ $=y_{c}+t_{y},$ $l^{\prime}$ $=l\cdot\exp(t_{l}),$ (3)
+$z_{c}^{\prime}$ $=z_{c}+t_{z},$ $h^{\prime}$ $=h\cdot\exp(t_{h}),$ (4)
+$\sin\theta^{\prime}$ $=\sin\theta+t_{\sin},$ $\cos\theta^{\prime}$ $=\cos\theta+t_{\cos}.$ (5)
+Once the box transformation has been applied, the center point and 8 corner points of the transformed box $\bm{b^{\prime}}$ are collected together as the anchor points $\{A^{i}\}_{i=0}^{8}$.
+Finally, the point-wise shift is independently applied to each anchor point $A^{i}$ to produce a PoI $P^{i}$ ($P^{i}=A^{i}+\Delta_{P}^{i}$).
 It’s worth noting that one object query corresponds to a set of PoIs derived from the center and corner points of the query box.
@@
 Since we exploit the voxel representation for point clouds, the location of the projected point on the BEV point cloud feature $F_{P}$ is computed as:
-{$m^{i}&=\frac{x^{i}-X_{\text{min}}}{(X_{\text{max}}-X_{\text{min}})\times d},\\ n^{j}&=\frac{y^{i}-Y_{\text{min}}}{(Y_{\text{max}}-Y_{\text{min}})\times d},\end{aligned}\right.$ (2)
+$\begin{cases}m^{i}&=\dfrac{x^{i}-X_{\text{min}}}{\left(X_{\text{max}}-X_{\text{min}}\right)\times d},\\[8.61108pt] n^{j}&=\dfrac{y^{i}-Y_{\text{min}}}{\left(Y_{\text{max}}-Y_{\text{min}}\right)\times d},\end{cases}\end{aligned}$ (6)
 where [$m^{i}$, $n^{i}$] is the projected location, [$X_{\text{min}}$, $Y_{\text{min}}$, $X_{\text{max}}$, $Y_{\text{max}}$] is the point cloud range in the BEV, and $d=8$ is the downsampling scalar of point cloud feature maps.
@@
 If the PoI is projected to only one view, this view will be leveraged to perform image feature sampling.
-Otherwise, we follow [5] to randomly choose one view for image feature sampling.
+Otherwise, we follow Cai et al.
+(2023) to randomly choose one view for image feature sampling.
 Given a 3D PoI $P^{i}$ in the LiDAR coordinate system, the projection onto the 2D image plane can be computed as follows:
@@
 where $P_{\text{image}}^{i}$ represents the coordinates of the projected PoI on the image plane, $\mathcal{I}$ is the intrinsic camera matrix, $\mathcal{R}$ is the rotation matrix of the extrinsic parameters, and $\mathcal{T}$ is the translation vector of the extrinsic parameters.
-In our camera encoder, we exploit an image backbone network followed by an FPN [35] to produce feature maps at P2, P3, P4, and P5 (1/4, 1/8, 1/16, and 1/32 downsampling, respectively).
+In our image encoder, we exploit an image backbone network followed by an FPN Lin et al.
+(2017a) to produce feature maps at P2, P3, P4, and P5 (1/4, 1/8, 1/16, and 1/32 downsampling, respectively).
 For a given projected point $P_{\text{image}}^{i}$, a set of image feature $\{\bm{g}^{i}_{j}\}_{j=2}^{5}$ is sampled from the multi-scale feature maps via bilinear interpolation.
@@
 $\bm{f}_{I}^{i}=\frac{\sum_{j=2}^{5}\text{exp}(w_{j}^{i})\cdot\bm{g}_{j}^{i}}{\sum_{j=2}^{5}\text{exp}(w_{j}^{i})},$ (8)
-where $\bm{f}_{I}^{i}$ represents the aggregated image feature and $\{w^{i}_{j}\}_{j=2}^{5}$ are scale weight coefficients, which are predicted by a linear layer according to the query feature $\bm{v}$.
-Figure 3: An illustration of our dynamic fusion block.
+where $\bm{f}_{I}^{i}$ represents the aggregated image feature and $\{w^{i}_{j}\}_{j=2}^{5}$ are scale weight coefficients, which are predicted by a linear layer with the query feature $\bm{v}$.
 3.5 Dynamic Multi-Modal Feature Fusion
@@
 Finally, the aggregated feature is added back to the query feature.
-The dynamic linear layers and the conventional linear layer leveraged for PoI feature aggregation are followed by a Layer Normalization operation [1] and a ReLU activation layer.
+The dynamic linear layers and the conventional linear layer leveraged for PoI feature aggregation are followed by a Layer Normalization operation Ba et al.
+(2016) and a ReLU activation layer.
+Figure 3: An illustration of our dynamic multi-modal feature fusion block, which first fuses image and point cloud feature at each PoI, and then integrates the PoIs of the same object query in canonical order.
 3.6 Prediction Head and Training Objective
@@
 In the classification head, we predict the binary classification score for each category.
-In the regression head, we follow the iterative refinement scheme [55, 84, 37] to predict the center delta over the predicted box center from the previous iteration.
+In the regression head, we follow the iterative refinement scheme Teed and Deng (2020); Zhu et al.
+(2020); Lin et al.
+(2023) to predict the center delta over the predicted box center from the previous iteration as follows:
+$\bm{c}_{t}=\bm{c}_{t-1}+\Delta\bm{c}_{t},$ (9)
+where $\bm{c}_{t}$ is the predicted box center predicted in the $t$-th recurrent iteration and $\Delta\bm{c}_{t}$ is the center delta.
 The other parts of the predicted 3D bounding box, including the box dimension, the heading direction, and the velocity, are independently predicted at each time.
 Training Objective.
-We follow the practice of set prediction [6] for target assignment.
-Specifically, we exploit bipartite matching to perform one-to-one assignments between the predicted 3D bounding boxes and the ground-truth boxes.
-In corresponding to the prediction head, our training objective also includes two parts: a focal loss [36] for the classification head and an L1 loss for the regression head.
-The overall training objective is computed as:
+We follow the practice of set prediction Carion et al.
+(2020) for the target assignment.
+Specifically, bipartite matching is exploited to obtain one-to-one assignments between predicted results and ground-truth ones.
+In corresponding to the prediction head, our training objective also includes two parts: a focal loss Lin et al.
+(2017b) for the classification head and an L1 loss for the regression head.
+The overall training objective is computed as follows:
 $L=\alpha\cdot L_{cls}(\bm{y}_{gt},\bm{y}_{pred})+\beta\cdot L_{reg}(\bm{b}_{gt},\bm{b}_{pred})$ (10)
-where $\bm{y}_{pred}$ is the classification score, $\bm{b}_{pred}$ is the predicted 3D box, $\alpha$ and $\beta$ are set as 2.0 and 0.25 following [69, 40] to balance the losses.
-The training objective is applied to the predicted output of each recurrent iteration.
+where $\bm{y}$ indicates classification logits, $\bm{b}$ means box coordinates, and $\alpha$ and $\beta$ are coefficients that balance these two losses.
+We set $\alpha$ to 2.0 and set $\beta$ to 0.25 following Yan et al.
+(2023); Liu et al.
+(2022).
 4 Experiments
-Method Present at Modality Camera Backbone LiDAR Backbone val test
-NDS $\uparrow$ mAP $\uparrow$ NDS $\uparrow$ mAP $\uparrow$
-DETR3D [63] CORL’21 C ResNet-50 - 43.4 34.9 47.9 41.2
-BEVFormer [33] ECCV’22 C ResNet-50 - 51.7 41.6 56.9 48.1
-CenterPoint [76] CVPR’21 L - VoxelNet 66.8 59.6 67.3 60.3
-TransFusion-L [2] CVPR’22 L - VoxelNet 70.1 65.1 70.2 65.5
-${}^{\star}$  [77] NeurIPS’21 C+L ResNet-50 VoxelNet 70.8 67.1 70.5 66.4
-PointAugmenting [58] CVPR’21 C+L ResNet-50 VoxelNet - - 71.0 66.8
-FUTR3D [8] CVPRW’22 C+L ResNet-101 VoxelNet 68.3 64.5 - -
-TransFusion [2] CVPR’22 C+L ResNet-50 VoxelNet 71.3 67.5 71.6 68.9
-UVTR [29] NeurIPS’22 C+L ResNet-101 VoxelNet 70.2 65.4 71.1 67.1
-BEVFusion (PKU) [34] NeurIPS’22 C+L Dual-Swin-T VoxelNet 72.1 69.6 73.3 71.3
-DeepInteraction [75] NeurIPS’22 C+L ResNet-50 VoxelNet 72.6 69.9 73.4 70.8
-BEVFusion (MIT) [42] ICRA’23 C+L Swin-T VoxelNet 71.4 68.5 72.9 70.2
-${}^{\star}$ [26] CVPR’23 C+L ResNet-50 VoxelNet 72.0 69.3 74.0 71.5
-${}^{\star}$ [31] arXiv’23 C+L ResNet-50 Sparse-ResUNet 72.7 70.4 74.0 70.6
-UniTR [60] ICCV’23 C+L Multi-Modal DSVT Multi-Modal DSVT 73.1 70.0 74.1 70.5
-SparseFusion [67] ICCV’23 C+L ResNet-50 VoxelNet 72.8 70.4 73.8 72.0
-${}^{\triangle}$ [69] ICCV’23 C+L VoVNet-99 VoxelNet 72.9 70.3 74.1 72.0
-ObjectFusion [5] ICCV’23 C+L Swin-T VoxelNet 72.3 69.8 73.3 71.0
-PoIFusion (ours) - C+L ResNet-50 VoxelNet 73.2 71.2 74.3 72.6
-PoIFusion (ours) - C+L Swin-T VoxelNet 73.6 71.7 74.9 73.4
-Table 1: Comparison of 3D object detection performance on nuScenes dataset.
-“C” means camera data (RGB images), “L” means LiDAR data (point clouds), $\star$” indicates the image input depends on predicted 2D instance masks, and $\triangle$” indicates the image resolution is enlarged to $1600\times 640$.
-We highlight the top-2 entries with bold font.
-We do not apply any test-time augmentation or model ensemble.
-4.1 Experimental Setup
-Dataset and Metric.
-We evaluate our method on the nuScenes dataset [4], which includes 700, 150, and 150 driving scenes for training, validation, and testing.
+Table 1: Performance comparison with state-of-the-art methods on nuScenes validation/test split.
+C: camera data (RGB images).
+L: LiDAR data (point clouds).
+$\star$: the image input depends on predicted 2D instance masks.
+$\triangle$: the image resolution of CMT is enlarged to $1600\times 640$.
+Our approach does not utilize test-time augmentation or employ model ensembling.
+Methods Modality Camera Backbone LiDAR Backbone val test
+mAP $\uparrow$ NDS $\uparrow$ mAP $\uparrow$ NDS $\uparrow$
+DETR3D Wang et al. (2022) C ResNet-50 - 34.9 43.4 41.2 47.9
+BEVFormer Li et al. (2022c) C ResNet-50 - 41.6 51.7 48.1 56.9
+CenterPoint Yin et al. (2021a) L - VoxelNet 59.6 66.8 60.3 67.3
+TransFusion-L Bai et al. (2022) L - VoxelNet 65.1 70.1 65.5 70.2
+MVP  Yin et al. (2021b) C+L ResNet-50 VoxelNet 67.1 70.8 66.4 70.5
+PointAugmenting Wang et al. (2021a) C+L ResNet-50 VoxelNet - - 66.8 71.0
+FUTR3D Chen et al. (2022a) C+L ResNet-101 VoxelNet 64.5 68.3 - -
+TransFusion Bai et al. (2022) C+L ResNet-50 VoxelNet 67.5 71.3 68.9 71.6
+UVTR Li et al. (2022a) C+L ResNet-101 VoxelNet 65.4 70.2 67.1 71.1
+BEVFusion (PKU) Liang et al. (2022) C+L Dual-Swin-T VoxelNet 69.6 72.1 71.3 73.3
+DeepInteraction Yang et al. (2022c) C+L ResNet-50 VoxelNet 69.9 72.6 70.8 73.4
+BEVFusion (MIT) Liu et al. (2023b) C+L Swin-T VoxelNet 68.5 71.4 70.2 72.9
+MSMDFusion Jiao et al. (2023) C+L ResNet-50 VoxelNet 69.3 72.0 71.5 74.0
+UniTR Wang et al. (2023b) C+L Multi-Modal DSVT Multi-Modal DSVT 70.0 73.1 70.5 74.1
+SparseFusion Xie et al. (2023) C+L ResNet-50 VoxelNet 70.4 72.8 72.0 73.8
+CMT Yan et al. (2023) C+L VoVNet-99 VoxelNet 70.3 72.9 72.0 74.1
+ObjectFusion Cai et al. (2023) C+L Swin-T VoxelNet 69.8 72.3 71.0 73.3
+FSF Li et al. (2024) C+L ResNet-50 Sparse-ResUNet 70.4 72.7 70.6 74.0
+PoIFusion (ours) C+L ResNet-50 VoxelNet 71.2 73.2 72.6 74.3
+PoIFusion (ours) C+L Swin-T VoxelNet 71.7 73.6 73.4 74.9
+4.1 Dataset and Evaluation Metric
+We evaluate our approach on nuScenes Caesar et al. (2020) and Argoverse2 Wilson et al. (2021) datasets.
+NuScenes.
+The nuScenes dataset includes 700, 150, and 150 driving scenes for training, validation, and testing, respectively.
 The vehicle for data collection is equipped with a 32-beam LiDAR sensor and 6 surrounding-view RGB cameras, thus providing both point clouds and multi-view images.
@@
 Both the multi-sensor data and annotation facilitate the exploration of multi-modal 3D object detection.
-We follow the official policy to mainly evaluate our method in terms of mean average precision (mAP) and nuScenes detection score (NDS).
+We follow the official policy to mainly evaluate our method on the 3D object detection benchmark in terms of mean average precision (mAP) and nuScenes detection score (NDS).
 The mAP is averaged over distance thresholds $0.5m$, $1m$, $2m$, and $4m$ on the BEV across $10$ classes.
 NDS is a weighted average of mAP and other true-positive metrics including mATE, mASE, mAOE, mAVE, and mAAE.
+Argoverse2.
+The Argoverse 2 (AV2) dataset is a large-scale benchmark for perception and prediction in autonomous driving.
+It comprises 150,000 annotated frames, which is five times larger than the nuScenes dataset, and 1,000 driving scenes.
+It features two 32-beam LiDARs combined into a 64-beam LiDAR and seven high-resolution surrounding cameras, offering a full 360° field of view and a valid detection range of up to 200 meters, covering an area of 400m × 400m.
+The dataset is particularly suited for long-range multi-modal object detection tasks.
+We evaluate our method across 26 categories.
+For evaluation, in addition to mean Average Precision (mAP), AV2 provides the Composite Detection Score (CDS), a comprehensive metric that combines mAP with other true positive metrics like Average Translation Error (mATE), Average Scale Error (mASE), and Average Orientation Error (mAOE).
+4.2 Experimental Setup
 Network Configuration.
-The camera encoder is composed of an image backbone network (i.e., ResNet [23] or Swin-Transformer [41]) and an FPN [35] to enrich the multi-scale information.
-The LiDAR encoder consists of a 3D sparse VoxelNet [80, 70] and a BEV backbone network.
-We follow the common practice of setting the image resolution as $800\times 448$ and setting the voxel size as ($0.075m$, $0.075m$, $0.2m$).
-In our multi-modal object decoder, there are $900$ object queries.
-The feature dimension of query embedding is set as $256$.
+The image encoder is composed of an image backbone network (i.e., ResNet He et al.
+(2016) or Swin-Transformer Liu et al.
+(2021)) and an FPN Lin et al.
+(2017a) to enrich the multi-scale information.
+The voxel encoder comprises of a 3D voxel backbone network and a BEV backbone network.
+For nuScenes, we follow the common practice of using VoxelNet Zhou and Tuzel (2018) as the 3D backbone network, setting the image resolution as $800\times 448$, and setting the voxel size as ($0.075m$, $0.075m$, $0.2m$).
+The number of object queries is set as $900$.
+For Argoverse2, there is less literature that can be referred to.
+Therefore, we follow FSF Li et al.
+(2024) to use SparseResUNet as the 3D backbone network.
+The image resolution is $960\times 640$ and the voxel size is ($0.2m$, $0.2m$, $0.2m$).
+It is worth noting that FSF keeps the output of SparseResUNet without downsampling, while we downsample the output for 4 times to decrease the resolution of BEV features.
+Since applied for a larger detection range, we use $1600$ object queries on Argoverse2.
+The configuration of our PoI multi-modal detector on both datasets keeps the same.
+Specifically, the feature dimension of query embedding is set as $256$.
 The channel of image and point cloud features is also transformed to $256$ before being fed into the multi-modal decoder.
 We equally divide the channels of feature maps of each modality into 4 groups, and generate PoIs for each group.
-This grouping operation improves the capacity of the network [66] and reduces the computation costs of the dynamic fusion block.
+This grouping operation improves the capacity of the network Xie et al.
+(2017) and reduces the computation costs of the dynamic fusion block.
 Moreover, the first dynamic fusion layer halves the channel of the query embedding, and the second dynamic fusion layer restores the feature dimension.
-The multi-modal object decoder is iteratively applied 6 times, with shared parameters.
+The PoI multi-modal decoder is iteratively applied 6 times, with shared parameters.
 Training and Inference.
-We follow the same training pipeline of previous methods [42, 5, 67, 34, 75], and use the same pretrained models.
-Specifically, the camera encoder is initialized with the model weights of a Mask R-CNN [24] model pre-trained on nuImage [4] dataset.
-The LiDAR encoder is initialized with the model weights of pretrained TransFusion-L [2].
-Our fusion framework is trained for 6 epochs with the AdamW optimizer.
-The training sample is resampled by CBGS [82] strategy.
+To make fair comparison, we follow the same training pipeline of previous methods Liu et al.
+(2023b); Cai et al.
+(2023); Xie et al.
+(2023); Liang et al.
+(2022); Yang et al.
+(2022c); Li et al.
+(2024), and use the same pretrained models.
+Specifically, the image encoder is pre-trained on nuImage Caesar et al.
+(2020) dataset.
+For nuScenes, the voxel encoder is initialized with the model weights of pretrained TransFusion-L Bai et al.
+(2022).
+For Argoverse2, the 3D voxel backbone network is initialized with the model weights of pretrained FSD Fan et al.
+(2022b), while the BEV backbone network is randomly initialized.
+For both datasets, our fusion framework is trained for 6 epochs with the AdamW optimizer.
+For nuScenes, the training sample is resampled by CBGS Zhu et al.
+(2019) strategy.
 The initial learning rate is set as 1e-4, adapted with the one-cycle learning rate policy, and the weight decay is set as 0.01.
-To speed up training convergence, the query denoising technique [39, 69, 38] is utilized.
 Data augmentation including random flip, random rotation, random translation, random scaling, and random modal masking is adopted.
-We exploit 8 GPUs for training, with 2 samples on each GPU.
+We exploit 8 GPUs for training, with 2 samples on each GPU for nuScenes and 1 sample on each GPU for Argoverse2.
 At inference, PoIFusion outputs the top 300 detection boxes without NMS. We don’t exploit any test time augmentation or model ensemble techniques.
-4.2 Comparison with Other Methods
-Main Results of Detection.
-In Table 1, we conduct a comprehensive comparison on the nuScenes 3D object detection benchmark.
-The methods are divided into three groups: camera-based (“C”), LiDAR-based (“L”), and multi-modal (“C+L”).
+Table 2: Performance comparison with state-of-the-art methods on Argoverse 2 validation split.
+C-Barrel: construction barrel.
+MPC-Sign: mobile pedestrian crossing sign.
+A-Bus: articulated bus.
+C-Cone: construction cone.
+V-Trailer: vehicular trailer.
+Some categories are excluded from the table due to the limited number of instances they contain.
+However, the average results consider all categories, even those that are omitted.
+In this experiment, the input voxel size of our PoIFusion is (0.2m, 0.2m, 0.2m), the image resolution is $960\times 640$, and the image backbone is ResNet-50, following the setting of FSF Li et al.
+(2024).
+Methods
+Average
+Vehicle
+Bus
+Pedestrian
+Box Truck
+C-Barrel
+Motorcyclist
+MPC-Sign
+Motorcycle
+Bicycle
+A-Bus
+School Bus
+Truck Cab
+C-Cone
+V-Trailer
+Bollard
+Sign
+Large Vehicle
+Stop Sign
+Stroller
+Bicyclist
+mAP Far3D Jiang et al. (2024) 24.4 - - - - - - - - - - - - - - - - - - - -
+CenterPoint Yin et al. (2021a) 22.0 67.6 38.9 46.5 40.1 32.2 28.6 27.4 33.4 24.5 8.7 25.8 22.6 29.5 22.4 37.4 6.3 3.9 16.9 0.5 20.1
+FSD Fan et al. (2022b) 28.2 68.1 40.9 59.0 38.5 42.6 39.7 26.2 49.0 38.6 20.4 30.5 14.8 41.2 26.9 41.8 11.9 5.9 29.0 13.8 33.4
+VoxelNeXt Chen et al. (2023a) 30.5 72.0 39.7 63.2 39.7 64.5 46.0 34.8 44.9 40.7 21.0 27.0 18.4 44.5 22.2 53.7 15.6 7.3 40.1 11.1 34.9
+FSF Li et al. (2024) 33.2 70.8 44.1 60.8 40.2 50.9 48.9 28.3 60.9 47.6 22.7 36.1 26.7 51.7 28.1 41.1 12.2 6.8 27.7 25.0 41.6
+CMT Yan et al. (2023) 36.1 71.9 41.5 61.2 38.4 62.2 58.2 40.3 55.1 40.3 24.7 47.3 23.6 59.5 23.8 49.6 28.4 6.8 54.7 4.6 50.2
+PoIFusion (ours) 40.6 77.6 49.4 70.6 40.7 75.0 67.1 44.0 66.9 55.1 26.9 49.2 32.5 64.6 31.9 59.0 28.1 8.0 44.0 11.4 55.1
+CDS Far3D Jiang et al. (2024) 18.1 - - - - - - - - - - - - - - - - - - - -
+CenterPoint Yin et al. (2021a) 17.6 57.2 32.0 35.7 31.0 25.6 22.2 19.1 28.2 19.6 6.8 22.5 17.4 22.4 17.2 28.9 4.8 3.0 13.2 0.4 16.7
+FSD Fan et al. (2022b) 22.7 57.7 34.2 47.5 31.7 34.4 32.3 18.0 41.4 32.0 15.9 26.1 11.0 30.7 20.5 30.9 9.5 4.4 23.4 11.5 28.0
+VoxelNeXt Chen et al. (2023a) 23.0 57.7 30.3 45.5 31.6 50.5 33.8 25.1 34.3 30.5 15.5 22.2 13.6 32.5 15.1 38.4 11.8 5.2 30.0 8.9 25.7
+FSF Li et al. (2024) 25.5 59.6 35.6 48.5 32.1 40.1 35.9 19.1 48.9 37.2 17.2 29.5 19.6 37.3 21.0 29.9 9.2 4.9 21.8 18.5 32.0
+CMT Yan et al. (2023) 27.8 62.2 33.6 46.8 30.8 47.3 47.6 30.2 43.1 29.8 18.9 38.4 16.9 42.5 17.1 34.5 21.1 5.0 43.0 3.2 40.4
+PoIFusion (ours) 31.6 66.5 40.8 54.8 33.0 58.4 54.6 28.7 55.0 42.8 20.4 40.0 24.7 47.8 23.5 42.3 21.6 5.8 35.0 8.7 42.8
+4.3 Comparison with State-of-the-art Methods
+4.3.1 Results on nuScenes
+In Table 1, we make a comprehensive comparison on nuScenes 3D object detection benchmark.
+The compared methods are divided into three groups: camera-based methods (“C”), LiDAR-based methods(“L”), and multi-modal ones(“C+L”).
+All the compared multi-modal 3D object detection approaches are without temporal information.
+Both the test time augmentation and the model ensembling techniques are not used in this comparison.
 Broadly speaking, the multi-modal methods outperform the methods that leverage only one single modality, validating the benefits of integrating the semantic-intensive RGB image and location-aware LiDAR point clouds.
-Method
-Modality
-AMOTA (%) $\uparrow$
-AMOTP (%) $\downarrow$
-CenterPoint [76]
-L
-63.7
-60.6
-VoxelNeXt [11]
-L
-70.2
-64.0
-TransFusion [2]
-C+L
-71.8
-60.3
-BEVFusion (MIT) [42]
-C+L
-72.8
-59.4
-ObjectFusion [5]
-C+L
-74.2
-54.3
-PoIFusion(ours)
-C+L
-75.1 50.7
-Table 2: Comparison of 3D multi-object tracking performance. The results are evaluated on the nuScenes validation set.
-Method mAP NDS Latency (ms)
-A100
-V100
-Transfusion [2]
-67.5
-71.3
-153.8
-190.3
-BEVFusion (MIT) [42]
-68.5
-71.4
-113.9 135.5
-DeepInteraction [75]
-69.9
-72.6
-204.1
-344.8
-CMT [69]
-70.3
-72.9
-210.8
-387.1
-${}^{\dagger}$ [69]
-70.3
-72.9
-159.7
-PoIFusion (ours)
-71.2 73.2
-115.2
-163.7
-Table 3: Runtime comparison.normal-$\dagger$” means this model is accelerated with the Flash-Attention Operator [15]. NVIDIA V100 GPU is not compatible with the Flash Attention package.
+With the same image resolution and the same voxel size, our proposed PoIFusion achieves state-of-the-art performance on the nuScenes dataset.
+Specifically, equipped with ResNet-50, our method achieves 73.2% NDS / 71.2% mAP on the nuScenes validation set.
+Upgrading the image backbone to a more powerful Swin-T Liu et al.
+(2021), the performance improves to 73.6% NDS and 71.7% mAP.
+We also submit the inference results to the official test server, and it reports that our PoIFusion with Swin-T image backbone achieves 74.9% NDS and 73.4% mAP, which makes an absolute improvement of 0.8% NDS and 1.4% mAP over the previous best results Yan et al.
+(2023).
+Compared to unified-view approaches Liang et al.
+(2022); Liu et al.
+(2023b); Li et al.
+(2022a), our PoIFusion method preserves modal-specific information more effectively, leading to a significant improvement over the representative BEVFusion model Liu et al.
+(2023b), with gains of 2.0% in NDS and 3.2% in mAP on the test set.
+Additionally, in contrast to recent work such as ObjectFusion Cai et al.
+(2023), which generates object-centric features by fusing regional features through RoI Pooling, our method leverages multi-modal fusion at adaptive Points of Interest (PoIs), resulting in an improvement of 1.6% in NDS and 2.4% in mAP.
+The use of PoIs enhances the flexibility of sampling locations and enables fine-grained feature fusion, contributing to the superior performance of our approach.
+4.3.2 Results on Argoverse2
+To study the performance of long-range 3D object detection, which is of significance to ensure safe driving, we further validate our approach on Argoverse2 dataset.
+The performance comparison is presented in Table 2.
+Compared to nuScenes, the long-tail category distribution of Argoverse2 emphasizes the importance of semantic-sensitive image features, while the precise localization of such a large detection range heavily relies on LiDAR sensor.
+Such facts make multi-modal fusion a promising choice for better 3D object detection.
+As shown in the table, our method consistently outperforms other competitors to deal with the challenge of long-range 3D object detection.
+Remarkably, the proposed PoIFusion sets a state-of-the-art record on this benchmark, achieving 40.6% mAP and 31.6% CDS.
+Note that the performance gap between our method and the strongest competitor CMT Yan et al.
+(2023) on Argoverse2 is more significant than that of nuScenes, further demonstrating the advantage of our PoI-based fusion to extract object-relevant features for addressing more challenge detection tasks.
+Table 3: Performance comparison of 3D multi-object tracking on nuScenes validation split in terms of AMOTA (%) and AMOTP (%).
+Methods AMOTA $\uparrow$ AMOTP $\downarrow$
+CenterPoint Yin et al. (2021a) 63.7 60.6
+VoxelNeXt Chen et al. (2023a) 70.2 64.0
+TransFusion Bai et al. (2022) 71.8 60.3
+BEVFusion (MIT) Liu et al. (2023b) 72.8 59.4
+ObjectFusion Cai et al. (2023) 74.2 54.3
+PoIFusion(ours) 75.1 50.7
+4.3.3 Extended Results for Multiple Object Tracking
+In addition to the main results of 3D object detection, we evaluate the generalizability of our method on the 3D multiple object tracking task of the nuScenes dataset.
+Specifically, we follow CenterPoint Yin et al.
+(2021a) to adopt the “tracking-by-detection” scheme that offline links the detection boxes into tracking tubelets and to evaluate the performance in terms of AMOTA and AMOTP.
+We would like to clarify that the compared methods, except VoxelNeXt Chen et al.
+(2023a), all follow the same “tracking-by-detection” scheme.
+As shown in Table 3, our PoIFusion works the best (75.1% AMOTA and 50.7% AMOTP) among the compared methods.
+This experiment shows that our better detection results can facilitate the downstream task, i.e., 3D multiple object tracking, which is also an important component in the perception system of an autonomous-driving vehicle.
+Table 4: Latency comparison.
+The latency is evaluated on nuScene, with batch size set as 1.
+$\dagger$”: accelerated with Flash-Attention Dao et al.
+(2022) (NVIDIA V100 GPU is not compatible with the Flash Attention operator).
+Top-2 entries are with bold font.
+Methods Latency (ms)
+A100 V100
+TransFusion Bai et al. (2022) 153.8 190.3
+BEVFusion (MIT) Liu et al. (2023b) 113.9 135.5
+DeepInteraction Yang et al. (2022c) 204.1 344.8
+CMT Yan et al. (2023) 210.8 387.1
+CMT Yan et al. (2023) 159.7 -
+PoIFusion (ours) 115.2 163.7
+4.3.4 Latency Comparison
+Furthermore, we conduct the latency comparison among the recent works.
+For a fair comparison, we re-evaluate the latency on an NVIDIA A100 and an NVIDIA V100 GPU.
+Since the implementation of the voxelization operation will not influence the detection results but becomes a confounder for runtime evaluation, we exclude the costed time of voxelization in this comparison.
+As shown in Table 4, the latency of our PoIFusion is 115.2ms on an NVIDIA A100 GPU, which is comparable to the fastest method, i.e., BEVFusion (MIT), which takes 113.9ms to process each sample.
+This comparison further demonstrates the potential of our proposed PoIFusion to serve as a strong baseline for future investigation and application.
+Table 5: Ablative experiments of components in the PoI multi-modal decoder.
+The backbone networks are ResNet-50 and VoxelNet.
+Image resolution is set as $800\times 448$, and the voxel size of point clouds is ($0.075m$,$0.075m$,$0.2m$).
+Our default setting is highlighted in gray.
+Anchor points NDS (%) mAP (%)
+Center only 72.8 70.6
+Center + corner 73.2 71.2
+(a) The choice of anchor points. Exploiting both the center and corner points as anchor points for points of interest generation yields better performance.
+PoI generation NDS (%) mAP (%)
+Baseline 72.4 70.1
++ B.T. 72.8 70.8
++ B.T. & P.S. 73.2 71.2
+(b) Operations to derive PoIs. The baseline in this experiment directly uses anchor points as PoIs. B.T.: box transformation. P.S.: point shift.
+Fusion block NDS (%) mAP (%)
+Static fusion 71.8 69.3
+Dynamic fusion 73.2 71.2
+(c) Fusion block. Dynamic fusion block produces adaptive parameters for the fusion layer, improving the performance.
+4.4 Experimental Analysis
+In this section, we conduct several groups of controlled experiments to experimentally analyze the effects of each factor in our proposed PoIFusion framework.
+In these experiments, the image backbone network is ResNet-50 and the point cloud backbone network is VoxelNet.
+Unless specified, the image resolution is set as $800\times 448$ and the voxel size is set as ($0.075m$,$0.075m$,$0.2m$).
+The models are all evaluated on the validation set of the nuScenes dataset.
+4.4.1 Components of PoI Multi-Modal Decoder
+In this section, we provide a detailed analysis of the key components in the PoI multi-modal decoder.
+The experiments are structured to investigate three critical aspects: (a) the selection of anchor points, (b) the generation of PoIs from anchor points, and (c) the fusion strategy to integrate multi-modal information.
+As shown in Table 5, the model that only uses the center point as the anchor to generate PoIs achieves 72.8% NDS and 70.6% mAP.
+By additionally involving the corner points as the anchor points, the NDS and mAP boosts to 73.2% and 71.2%, respectively.
+This comparison shows the benefits of representing the geometric property of query boxes.
+In Table 5, we validate the operations to derive PoIs from anchor points.
+The baseline directly takes the anchor points as PoIs without adaptation, achieving 70.1% mAP.
+By involving the holistic box-level transformation and further applying the point-wise shift, the performance boosts to 70.8 % mAP and 71.2% mAP, respectively.
+The online adjustment of PoIs ameliorates the misalignment in feature sampling and thereby boosts the detection result.
+Table 5 compares the effect of different fusion schemes, i.e., static fusion versus dynamic fusion.
+In static fusion, two conventional linear layers are exploited to process the concatenated multi-modal features at each PoI.
+The dynamic fusion scheme refers to our proposed on in Section 3.5.
+Thanks to the adjustment of fusion layers’ parameters for each object query, our dynamic fusion outperforms the static setting by 1.4% NDS and 1.9% mAP.
+Table 6: Effect of leveraging different modalities. Integrating RGB images and LiDAR point clouds remarkably boosts the detection result.
+Modality NDS (%) mAP (%)
+Camera only 40.6 29.4
+LiDAR only 70.2 65.3
+Camera + LiDAR 73.2 71.2
+4.4.2 Effect of Different Modalities
+In this experiment, we train models with single-modality input following the same training setting as introduced in Section 4.2.
+This experiment is presented in Table 6.
+The image-only baseline achieves 40.6% NDS and 29.4% mAP, and the LiDAR-only baseline achieves 70.2% NDS and 65.3% mAP.
+Integrating the multi-modal information remarkably boosts the performance to $73.2\%$ NDS and $71.2\%$ mAP.
+Please note that the performance of our LiDAR-only baseline is comparable to that of TransFusion-L (70.1% NDS and 65.1% mAP), which is the LiDAR-only baseline of Bai et al.
+(2022); Liang et al.
+(2022); Liu et al.
+(2023b); Yang et al.
+(2022c); Xie et al.
+(2023); Cai et al.
+(2023), but our fusion model makes a more significant improvement.
+This comparison further validates the benefits of our proposed fusion at points of interest scheme to integrate complementary information from images and point clouds.
+4.4.3 Effect of the Input Resolution
+The input resolution (i.e., image resolution, and voxel size) is also one of the key factors that influence the detection result.
+In this experiment, we present the performance comparison under various input resolutions.
+Image Resolution.
+As shown in Table 7, we enlarge the image resolution from $800\times 448$ to $1600\times 640$, which is the same as the input resolution of the strongest competitor CMT.
+Compared to our default setting, the detection performance is observed with an improvement of 0.5% NDS and 0.7% mAP.
+The larger resolution provides more detailed textual information from the RGB images and thus benefits multi-modal 3D object detection.
+It is worth noting that although increase the image resolution can definitely improve the performance, we have not increase it in our default setting, which ensures fair comparison with others.
+Table 7: Ablative experiments of the input resolution. The backbone networks are ResNet-50 and VoxelNet. Our default setting is highlighted in gray.
+Image resolution NDS (%) mAP (%)
+$800\times 448$ 73.2 71.2
+$1600\times 640$ 73.7 71.9
+(a) Image Resolution. Increasing the image resolution strengthens the textual information provided by images.
+Voxel size NDS (%) mAP (%)
+(0.125, 0.125, 0.2) 71.6 69.1
+(0.1, 0.1, 0.2) 72.3 69.8
+(0.075, 0.075, 0.2) 73.2 71.2
+(b) Voxel Size. A smaller voxel size corresponds to a larger resolution of the point cloud feature, facilitating the localization information for 3D object detection.
+Voxel Size.
+In Table 7, we further analyze the influence of the voxel size by setting the voxel size as (0.125m, 0.125m, 0.2m), (0.1m, 0.1m, 0.2m) and (0.075m, 0.075m, 0.2m).
+As shown in the table, the model with a smaller voxel size works better than that with a larger one.
+Smaller voxel size results in a larger resolution of the point cloud feature map, which facilitates more precise localization information.
 Figure 4: Qualitative results on the nuScenes validation set.
@@
 We use different colors for different categories.
-NDS
-mAP
-center only
-72.8
-70.6
-center + corner
-73.2 71.2
-NDS
-mAP
-baseline
-72.4
-70.1
-+ B.T.
-72.8
-70.8
-+ B.T. & P.S.
-73.2 71.2
-NDS
-mAP
-static fusion
-71.8
-69.3
-dynamic fusion
-73.2 71.2
-NDS
-mAP
-camera only
-40.6
-29.4
-LiDAR only
-70.2
-65.3
-camera + LiDAR
-73.2 71.2
-image resolution
-NDS
-mAP
-800$\times$448
-73.2
-71.2
-1600$\times$640
-73.7 71.9
-voxel size
-NDS
-mAP
-(0.125, 0.125, 0.2)
-71.6
-69.1
-(0.1, 0.1, 0.2)
-72.3
-69.8
-(0.075, 0.075, 0.2)
-73.2 71.2
-Table 4: Ablative experiments.
-In these experiments, the backbone networks are ResNet-50 and VoxelNet.
-Unless specified, image resolution is set as $800\times 448$, and the voxel size of point clouds is ($0.075m$,$0.075m$,$0.2m$).
-Our default setting is highlighted in gray.
-Our proposed PoIFusion has set a state-of-the-art record on the nuScenes dataset.
-Specifically, equipped with ResNet-50, our method achieves 73.2% NDS / 71.2% mAP on the validation set and achieves 74.3% NDS / 72.6% mAP on the test set.
-Upgrading the image backbone to a more powerful Swin-T [41], the performance on the test set boosts to 74.9% NDS and 73.4% mAP, which makes an absolute improvement of 0.8% NDS and 1.4% mAP over the previous best results [69].
-Compared to the unified-view methods [34, 42, 29], our PoIFusion better keeps modal-specific information and thus outperforms the representative BEVFusion [42] by 2.0% NDS and 3.2% mAP on the nuScenes test set.
-In comparison with recent work, i.e., ObjectFusion [5], which also generates object-centric features, our method performs multi-modal fusion at adaptive PoIs, instead of fusing the regional features extracted from different representation spaces.
-The employment of PoIs improves the flexibility of sampling location and enables fine-grained feature fusion, thus working better.
-Moreover, the qualitative results are shown in Figure 4, with one example from a sunny day and another example from a rainy night.
-Our PoIFusion can robustly perform 3D object detection under varying weather and lighting conditions.
-Extended Results of MOT.
-In addition to the main results of 3D object detection, we evaluate the generalizability of our method on the 3D multiple object tracking task of the nuScenes dataset.
-Specifically, we follow CenterPoint [76] to adopt the “tracking-by-detection” scheme that offline links the detection boxes into tracking tubelets.
-Note that other compared methods, except VoxelNeXt [11], also follow this same scheme.
-As shown in Table 2, our PoIFusion works the best (75.1% AMOTA and 50.7%AMOTP) among the compared methods, demonstrating our better detection results can also benefit other downstream tasks.
-Runtime Comparison.
-Furthermore, we conduct runtime comparisons among the recent works.
-For a fair comparison, we re-evaluate the latency on an NVIDIA A100 and an NVIDIA V100 GPU.
-Since the implementation of the voxelization operation is a confounder for runtime evaluation but will not influence the detection result, we exclude its time cost in comparison.
-As shown in Table 3, the latency of our PoIFusion is comparable to the fastest method, i.e., BEVFusion (MIT), which further shows the potential of our method to serve as a strong baseline for future investigation.
-4.3 Ablative Experiments
-Effect of Components in the Multi-Modal Decoder.
-In this section, we analyze the design of the proposed multi-modal decoder from three aspects: (a) anchor points, (b) operations to generate PoIs, and (c) the fusion block.
-As shown in Table LABEL:tab:ablation:anchor, the model that only uses the center point achieves 72.8% NDS and 70.6% mAP.
-By additionally involving the corner points, the NDS and mAP boosts to 73.2% and 71.2%, respectively.
-This comparison shows the benefits of representing the geometric property of query boxes.
-In Table LABEL:tab:ablation:transformation, we validate the operations to derive PoIs from anchor points.
-The baseline directly takes the anchor points as PoIs without adaptation, achieving 70.1% mAP.
-By involving the holistic box-level transformation and further applying the point-wise shift, the performance boosts to 70.8 % mAP and 71.2% mAP, respectively.
-The online adjustment of PoIs ameliorates the misalignment in feature sampling and thereby boosts the detection result.
-Table LABEL:tab:ablation:fusion_block compares the effect of different fusion schemes, i.e., static fusion versus dynamic fusion.
-In static fusion, two conventional linear layers are exploited to process the concatenated multi-modal features at each PoI.
-The dynamic fusion scheme refers to our proposed on in Section 3.4.
-Thanks to the adjustment of fusion layers’ parameters for each individual object query, our dynamic fusion outperforms the static setting by 1.4% NDS and 1.9% mAP.
-Effect of Different Modalities.
-In this experiment, we train models with single-modality input following the same setting as our default one.
-As shown in Table LABEL:tab:ablation:modality, the image-only baseline achieves 40.6% NDS and 29.4% mAP.
-Meanwhile, the LiDAR-only baseline achieves 70.2% NDS and 65.3% mAP, which is similar to TransFusion-L [2] (70.1% NDS and 65.1% mAP).
-Integrating the multi-modal information remarkably boosts the performance to $73.2\%$ NDS and $71.2\%$ mAP, validating the capability of our proposed fusion paradigm to integrate the complementary information from images and point clouds.
-Effect of the Input Resolution.
-The input resolution is also one of the key factors that influence the detection result.
-In Table LABEL:tab:ablation:img_resolution, we enlarge the image resolution from $800\times 448$ to $1600\times 640$, obtaining 0.5% NDS and 0.7% mAP gain.
-In Table LABEL:tab:ablation:voxel_size, we further analyze the influence of the voxel size.
-Either a larger image resolution or a smaller voxel size facilitates the feature representation of the corresponding modality, thus boosting the detection result.
-Our default setting employs an image resolution of $800\times 448$ and a voxel size of ($0.0075m,0.0075m,0.2m$), which is the same as most of recent works [2, 5, 67, 75].
-Offsets (m)
-0.0
-0.2
-0.4
-0.6
-0.8
-1.0
-NDS (%)
-73.2
-73.0
-72.9
-72.7
-72.3
-72.0
-mAP (%)
-71.2
-71.1
-70.8
-70.5
-70.0
-69.4
-Table 5: Sensor misalignment. Performance comparison on the impact of translation offsets due to calibration errors.
-4.4 Detection Robustness Analysis
-Sensor Misalignment.
-In real applications, sensors can be misaligned due to physical impacts, installation errors, time-related drift, or some other unexpected reasons.
-This experiment investigates the robustness of PoIFusion against sensor misalignment.
-Specifically, we follow [2] to randomly add translation offsets to the calibration matrix at inference.
-The errors are formulated as uniformly distributed noise, with the maximum offset value varying from $0.2m$ to $1.0m$.
-As shown in Table 5, even when the maximum translation offset is $1.0m$, our method can achieve 72.0% NDS and 69.4% mAP, demonstrating the robustness of our fusion paradigm against sensor misalignment.
-Sensor Failure.
-In addition to sensor misalignment, we further explore a more serious condition, i.e., sensor failure.
-To validate our method under camera failure cases, we randomly drop several images by filling the corresponding images as all zeros.
-The results with different numbers of dropped images are presented in Table 6.
-The performance keeps dropping along with the increasing number of cameras that fail to work.
-Even with half of the cameras blocked, our PoIFusion achieves 68.6% mAP, obviously better than the LiDAR-only method.
-Moreover, in the extreme case that all of the cameras are not working, our method works comparable to the LiDAR-only baseline.
-# Dropped Images
-0
-1
-3
-6
-NDS (%)
-73.2
-72.8
-72.0
-70.2
-mAP (%)
-71.2
-70.3
-68.6
-65.6
-Table 6: Camera failure. Performance comparison on the impact of dropped images due to camera failure.
-Discarded Sectors
-${}^{\circ}$
-${}^{\circ}$
-${}^{\circ}$
-${}^{\circ}$
-${}^{\circ}$
-NDS (%)
-73.2
-72.7
-72.0
-71.3
-70.7
-mAP (%)
-71.2
-70.2
-68.9
-67.6
-66.4
-Table 7: LiDAR failure. Performance comparison on the impact of discarded point clouds within a sector due to LiDAR failure.
-Furthermore, to simulate the failure of a LiDAR sensor, we randomly discard point cloud data within a specified sectorial region based on the azimuthal angle.
-As shown in Table 7, in general, the impact of LiDAR failure is more pronounced than that of camera failure, since the spatial information captured by the LiDAR sensor is essential for precise localization.
-Notwithstanding the total absence of point cloud data within a $24^{\circ}$ sector, the mAP achieved by our method is still 1.1% higher than the LiDAR-only baseline.
-These two experiments demonstrate the robustness of our proposed PoIFusion against Sensor Failure.
-5 Conclusion
-In this work, we present PoIFusion, a query-based multi-modal 3D object detector, which is characterized by performing multi-modal feature sampling and fusion at generated points of interest.
-On the one hand, PoIFusion keeps the original view of the image and point cloud feature to preserve the modal-specific information.
-On the other hand, the employment of points of interest facilitates fine-grained fusion and enables representing object features in a flexible way.
-Our PoIFusion sets a state-of-the-art record on the highly competitive nuScenes dataset, and in the meantime keeps efficient runtime.
-Moreover, the experimental analysis demonstrates our method is robust to sensor misalignment and failure.
-We suppose our PoIFusion would serve as a strong baseline with huge potential for future exploration.
-Supplementary Material
-This supplementary material is composed of three sections.
-In Section A, we introduce supplementary details of our implementation.
-Then, in Section B, we provide additional experiments to further validate the merits of our proposed method.
-Finally, in section C, we visualize more examples to help intuitively present the detection results.
-Appendix A Supplementary Details
-Query Initialization.
-As introduced in our main paper, we formulate each object query as a learnable dynamic 3D box $\bm{b}\in\mathbb{R}^{8}$ and an attached feature vector $\bm{v}\in\mathbb{R}^{256}$.
-Specifically, the query box is represented as:
-$\bm{b}=[x_{c},y_{c},z_{c},w,l,h,sin\theta,cos\theta],$ (6)
-where [$x_{c}$, $y_{c}$, $z_{c}$] is the center location, [$w$, $l$, $h$] is the box dimension, and $\theta$ is the azimuth angle.
-Notably, while velocity is a component of the detection results, it is not factored into the query boxes.
-When we initialize the query box, the BEV center location [$x_{c}$, $y_{c}$] is set to be uniformly distributed on the BEV, and the height center $z_{c}$ is set as 0.
-Besides, the dimensions $w$, $l$, $h$ of each box are respectively initialized as 6, 3, 2, and the azimuth is initialized as 0.
-Distance-Biased Self-Attention Layer.
-As outlined in Section 3.1, the multi-modal decoder utilizes a distance-biased self-attention layer [38] to capture the relationships and dependencies between different object queries.
-Here, we provide a comprehensive review of it.
-Specifically, a conventional self-attention layer [56] calculates the attention score based on the dot product of query and key, which is detailed as:
-[…]
-where Q and K correspond to the query and key, and C is the channel dimension.
-The distance-biased self-attention layer incorporates a bias term that is contingent upon the centroid distance between pairs of query boxes.
-The adjusted attention score is computed as:
-[…]
-where $D$ indicates the centroid distance between query boxes, and $\psi$ is a scalar produced by a linear layer according to the query feature.
-In our multi-modal object decoder, we use 8 heads in the self-attention layer.
-The $\psi$ is not shared in different heads.
-We emphasize that the distance-biased self-attention layer is not the contribution of our work.
-We implement the self-attention layer with it since it improves the adaptability of the model with negligible costs.
-Appendix B Additional Experimental Analysis
-Different Weathers and Lighting Conditions.
-The 3D object detection system on autonomous driving vehicles is supposed to be capable of working under varying lighting and weather conditions.
-In the main manuscript, we have shown two qualitative examples to give a glimpse of the robustness of our PoIFusion.
-In this experiment, we further follow [42] to divide the validation set of the nuScenes dataset into two pairs of subsets: sunny/rainy and day/night for comprehensive quantitative evaluation.
-The evaluated performance in terms of mAP (%) is summarized in Table S1.
-In general, the multi-modal 3D object detectors are not likely to be affected by the rainy weather.
-The changing of lighting condition significantly affects the detection performance, i.e., the mAP on the night subset is much lower than that of the day subset.
-Among the compared methods, our PoIFusion achieves the best performance under different weather and lighting conditions, further demonstrating the robustness of our fusion strategy.
+Table 8: Performance comparison between using shared parameters and unshared parameters for the PoI multi-modal decoder. Our default setting is highlighted in gray.
+Parameters NDS (%) mAP (%)
+Unshared parameters 73.3 70.9
+Shared parameters 73.2 71.2
+4.4.4 Shared Parameters VS. Unshared Parameters for the PoI Multi-Modal Decoder
+Our PoI multi-modal decoder is applied iteratively, with the option to use either shared or unshared parameters.
+The performance comparison between these two configurations is shown in Table 8, where both models demonstrate comparable results.
+However, employing shared parameters reduces the overall model size, making it more efficient.
+Consequently, we adopt the shared-parameter configuration as the default setting for the PoI multi-modal decoder.
+Table 9: Different weathers and lighting conditions.
+The reported performance is compared in terms of mAP (%) on the nuScenes validation set.
+We report the result of our PoIFusion with the Swin-T image backbone network, which is consistent with MVP Yin et al.
+(2021b) and BEVFusion Liu et al.
+(2023b).
 mAP (%) Modality Sunny Rainy Day Night
@@
 PoIFusion (ours) C+L 71.1 70.2 71.3 45.0
-Table S1: Different lighting and weather conditions.
+4.5 Robustness Analysis
+Besides the accuracy and efficiency, the capability of working under different environments and corruptions is also what we expected for a desirable perception module in autonomous driving vehicles.
+In this section, we present a series of experiments to examine the robustness of our PoIFusion framework.
+4.5.1 Different Weathers and Lighting Conditions
+The 3D object detection system on autonomous driving vehicles is supposed to be capable of working under varying lighting and weather conditions.
+In this experiment, we follow Liu et al.
+(2023b) to divide the validation set of the nuScenes dataset into two pairs of subsets: sunny/rainy and day/night for comprehensive quantitative evaluation.
+The evaluated performance in terms of mAP (%) is summarized in Table 9.
+In general, the multi-modal 3D object detectors are not likely to be affected by the rainy weather.
+The changing of lighting condition significantly affects the detection performance, i.e., the mAP on the night subset is much lower than that of the day subset.
+Among the compared methods, our PoIFusion achieves the best performance under different weather and lighting conditions, validating the robustness of fusion at points of interest.
+Moreover, we visualize the qualitative results of detecting 3D objects under different weather and lighting conditions in Figure 4.
+We present one example from a sunny day and another example from a rainy night.
+These examples further give an intuitive understanding that PoIFusion can robustly perform 3D object detection under varying weather and lighting conditions
+Table 10: Sensor misalignment.
+Performance comparison on the impact of translation offsets due to calibration errors.
 The reported performance is compared in terms of mAP (%) on the nuScenes validation set.
-We report the result of our PoIFusion with the Swin-T image backbone network.
-Effect of Iteration Times of the Multi-Modal Decoder.
-As introduced in Section 3.1, our multi-modal object decoder is iteratively applied to perform feature fusion and box refinement.
-In Table S2, we analyze how the iteration times of our multi-modal decoder influence the 3D object detection performance.
-Specifically, we vary the iteration times from 1 to 8 times and re-train the model for evaluation.
-In the beginning, the performance consistently improves along with the increasing iteration time.
-The peak performance is achieved when we apply the multi-modal object decoder 6 times.
-When we further increase the iteration time to 8, the performance drops for 0.5% NDS and 0.9% mAP.
-Therefore, in our default configuration, the multi-modal object decoder is iteratively applied 6 times.
-Moreover, as shown in the table, our method provides the potential to make a good trade-off between accuracy and latency.
-Specifically, when the iteration time is reduced to 4, our PoIFusion can still achieve the highest NDS (73.1%) and mAP (70.7%) among the compared methods presented in Table 1, and in the meantime run faster than the fastest competitor (i.e., BEVFusion) presented in Table 3.
-# Iteration Times
-NDS (%)
-mAP (%)
-Latency (ms)
-1
-67.1
-63.4
-87.7
-2
-71.9
-69.3
-94.2
-4
-73.1
-70.7
-99.8
-6
-73.2 71.2
-115.2
-8
-72.7
-70.3
-119.2
-Table S2: Effect of iteration times of the multi-modal decoder.
-The reported performance is compared on the nuScenes validation set.
-We exploit ResNet-50 as the image backbone network.
-The latency is evaluated on an NVIDIA A100 GPU.
-Shared Parameters VS. Unshared Parameters
-of the Multi-Modal Decoder.
-Since the multi-modal object decoder is iteratively applied multiple times, we can exploit either shared parameters or unshared parameters.
-The results of using shared parameters and unshared parameters are presented in Table S3.
-As shown in this comparison, these two models achieve comparable performance.
-Moreover, sharing parameters of the multi-modal object decoder makes a smaller model size.
-Therefore, we use shared parameters of the multi-modal decoder by default.
-NDS (%)
-mAP (%)
-Shared Parameters
-73.2
-71.2
-Unshared Parameters
-73.3
-70.9
-Table S3: Shared versus unshared parameters of the multi-modal object decoder. The models are evaluated on the nuScenes validation set, with ResNet-50 as the image backbone network.
-Training without the pretrained LiDAR encoder.
-In our main experiment, we follow the training policy of most recent works [2, 42, 67, 5, 75], i.e., starting with the model weight of pretrained TransFusion-L and training the fusion model for 6 epochs.
-Here, we further study whether our PoIFusion can also be trained without a pretrained TransFusion-L model.
-Specifically, in this experiment, we follow the training policy of CMT [69] that trains the fusion framework with 20 epochs, with only the pretrained image backbone network.
-The performance comparison between our PoIFusion and CMT is presented in Table S4.
-Please note that these two models are trained with the same training policy, backbone network, and input resolution (details are summarized in the caption of this table).
-As shown in the table, our PoIFusion achieves 72.3% NDS and 68.4% mAP without pretrained LiDAR encoder, outperforming CMT with the same input resolution and training policy by 1.5% NDS and 0.5% mAP.
-This experiment demonstrates our PoIFusion works very well under various training policies, further exhibiting its potential to serve as a strong baseline for future research.
-Method
-NDS (%)
-mAP (%)
-CMT
-70.8
-67.9
-PoIFusion (ours)
-72.3 68.4
-Table S4: Training without a pretrained TransFusion-L model.
-The compared models are trained with the same training policy and evaluated on the nuScenes validation set.
-Both models use ResNet-50 and VoxelNet as the backbone network, with image resolution set as $800\times 320$ and voxel size set as $(0.1m,0.1m,0.2m)$.
-Appendix C Additional Qualitative Results
-In our main manuscript, we only visualize two scenes due to the limitation of pages.
-Here, we add more qualitative examples in Figure S1 to help intuitively present the detection results of our method.
-From the visualized results, it can be seen that our proposed method is able to accurately detect 3D objects under complex road conditions.
-Figure S1: Supplement of Qualitative results on the nuScenes validation set.
+We report the result of our PoIFusion with the ResNet-50 image backbone network.
+Misalignment offset (m) NDS (%) mAP (%)
+0.0 73.2 71.2
+0.2 73.0 71.1
+0.4 72.9 70.8
+0.6 72.7 70.5
+0.8 72.3 70.0
+1.0 72.0 69.4
+4.5.2 Robustness Against Sensor Misalignment
+In real applications, sensors can be misaligned due to physical impacts, installation errors, time-related drift, or some other unexpected reasons.
+This experiment investigates the robustness of PoIFusion against sensor misalignment.
+Specifically, we follow Bai et al.
+(2022) to randomly add translation offsets to the calibration matrix at inference.
+The errors are formulated as uniformly distributed noise, with the maximum offset value varying from $0.2m$ to $1.0m$.
+As shown in Table 10, even when the maximum translation offset is $1.0m$, our method can achieve 72.0% NDS and 69.4% mAP, still improving our LiDAR-only baseline with 1.8% NDS and 4.1% mAP.
+This experiment demonstrates the robustness of our fusion paradigm against sensor misalignment.
+4.5.3 Robustness Against Sensor Failure
+In addition to sensor misalignment, we further explore a more terrible corruption, i.e., sensor failure.
+Camera Failure.
+To validate our method under camera failure cases, we randomly drop several images by filling the corresponding images as all zeros.
+The results with different numbers of dropped images are presented in Table 11.
+The performance keeps dropping along with the increasing number of cameras that fail to work.
+Even with half of the cameras blocked, our PoIFusion achieves 68.6% mAP, obviously better than the LiDAR-only method.
+Moreover, in the extreme case that all of the cameras are not working, our method still works a little bit better than the LiDAR-only baseline, showing the benefit of multi-modal joint training.
+Table 11: Camera failure. Performance comparison on the impact of dropped images due to camera failure.
+# Dropped Images NDS (%) mAP (%)
+0 73.2 71.2
+1 72.8 70.3
+3 72.0 68.6
+6 70.2 65.6
+Table 12: LiDAR failure. Performance comparison on the impact of discarded point clouds within a sector due to LiDAR failure.
+Discarded Sectors NDS (%) mAP (%)
+$0^{\circ}$ 73.2 71.2
+$6^{\circ}$ 72.7 70.2
+$12^{\circ}$ 72.0 68.9
+$18^{\circ}$ 71.3 67.6
+$24^{\circ}$ 70.7 66.4
+LiDAR Failure.
+Furthermore, to simulate the failure of a LiDAR sensor, we randomly discard point cloud data within a specified sectorial region based on the azimuthal angle.
+As shown in Table 12, in general, the impact of LiDAR failure is more pronounced than that of camera failure, since the information captured by the LiDAR sensor is essential for precise localization.
+Notwithstanding the total absence of point cloud data within a $24^{\circ}$ sector, the mAP achieved by our method is still 1.1% higher than the LiDAR-only baseline.
+These two experiments demonstrate the robustness of our proposed PoIFusion against Sensor Failure.
+5 Conclusion
+In this work, we introduce PoIFusion, a query-based multi-modal 3D object detector that leverages multi-modal feature sampling and fusion at strategically generated points of interest.
+PoIFusion preserves modal-specific information by maintaining the original views of both image and point cloud features, while the use of points of interest allows for fine-grained fusion and a flexible representation of object features.
+Our approach establishes a new state-of-the-art on the challenging nuScenes dataset, all while maintaining efficient runtime performance.
+Furthermore, experimental results highlight the robustness of PoIFusion to sensor misalignment and failure.
+We believe that PoIFusion will serve as a robust baseline with significant potential for future research and deployment in the field.
+6 Data Availability
+The nuScenes dataset Caesar et al.
+(2020) and Argoverse2 dataset Wilson et al.
+(2021) used in this study are well-recognized benchmarks which are public available.
+Specifically, nuScenes dataset can be accessed through https://www.nuscenes.org/ and Argoverse2 dataset can be accessed through https://www.argoverse.org/av2.html.
+We have not used extra private data in our experiments.